### v2

# The Mixed Integer Trust Region Problem

Abstract
In this paper we consider the problem of minimizing a general quadratic function over the mixed integer points in an ellipsoid. This problem is strongly NP-hard, NP-hard to approximate within a constant factor, and optimal solutions can be irrational. In our main result we show that an arbitrarily good solution can be found in polynomial time, if we fix the number of integer variables. This algorithm provides a natural extension to the mixed integer setting, of the polynomial solvability of the trust region problem proven by Ye, Karmarkar, Vavasis, and Zippel. As a result, our findings pave the way for designing efficient trust region methods for mixed integer nonlinear optimization problems. The techniques that we introduce are of independent interest and can be used in other mixed integer nonlinear optimization problems. As an example, we consider the problem of minimizing a general quadratic function over the mixed integer points in a polyhedron. For this problem, we show that a solution satisfying weak bounds with respect to optimality can be computed in polynomial time, provided that the number of integer variables is fixed. It is well known that finding a solution satisfying stronger bounds cannot be done in polynomial time, unless P=NP.
Key words: trust region problem; mixed integer quadratic programming; approximation algorithm; polynomial time algorithm; ellipsoid constraint
1 Introduction
Trust region methods form an important class of algorithms for nonlinear optimization problems. The appeal of these methods lies in their general applicability, the elegant theory, and their practical success. We refer the reader to the book [5] for a thorough introduction. In trust region methods, a quadratic function is used to model the nonlinear objective function, and this model is expected to be accurate (the model is “trusted”) only in a neighborhood of the current iterate $c$. To obtain the next iterate, first an affine change of variables is preformed to map $c$ to the origin, and then the model quadratic function is minimized over a neighborhood of the origin. If the 2-norm is used to define the neighborhood, the problem takes the form
$\begin{split}\min&\quad x^{\mathsf{T}}Hx+h^{\mathsf{T}}x\\ \operatorname{s.t.}&\quad x^{\mathsf{T}}x\leq 1.\end{split}$ (TR)
Here $H$ is a symmetric matrix in $\mathbb{Q}^{n\times n}$ and $h\in\mathbb{Q}^{n}$. Equation TR is often referred to, in the literature, as the trust region (sub)problem.
One key reason behind the success of trust region methods is the fact that Equation TR can be solved efficiently. In fact, while optimal solutions of Equation TR can be irrational (see Observation 3), an arbitrarily good solution can be found in polynomial time. This celebrated result is due to Ye [27], Karmarkar [14], and Vavasis and Zippel [26], and a formal statement is given in Lemma 3. Tractability results for extensions of Equation TR that include additional linear or quadratic constraints can be found, for example, in [12, 28, 3, 2, 6].
Trust region methods have been introduced also for mixed integer nonlinear optimization problems (see, e.g., [11, 10, 19, 20]). For this more general type of optimization problems, if the 2-norm is used to define the neighborhood, the next iterate can be obtained by solving the mixed integer trust region (sub)problem
$\begin{split}\min&\quad x^{\mathsf{T}}Hx+h^{\mathsf{T}}x\\ \operatorname{s.t.}&\quad x^{\mathsf{T}}x\leq 1\\ &\quad x\in\Pi_{p}(b^{1},\dots,b^{n})+\{c\}.\end{split}$ (MITR)
Here $H$ is a symmetric matrix in $\mathbb{Q}^{n\times n}$, $h,c\in\mathbb{Q}^{n}$, $b^{1},\dots,b^{n}\in\mathbb{Q}^{n}$ are linearly independent, $p\in\{0,1,\dots,n\}$, and $\Pi_{p}(b^{1},\dots,b^{n})$ denotes the mixed integer lattice
$\Pi_{p}(b^{1},\dots,b^{n}):=\left\{\sum_{i=1}^{n}\mu_{i}b^{i}:\mu_{i}\in\mathbb{Z}\ \forall i=1,\dots,p,\ \mu_{i}\in\mathbb{R}\ \forall i=p+1,\dots,n\right\},$
which is the image of the set $\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$ under an affine change of variables. As opposed to Equation TR, we will see in Section 3 that Equation MITR is strongly NP-hard and NP-hard to approximate within a constant factor, even in the pure integer case $p=n$. Even deciding whether the feasible region of Equation MITR is nonempty, is NP-hard.
The relationship between the computational complexity of Equations TR and MITR resembles the situation in linear optimization, where linear programs can be solved in polynomial time, while mixed integer linear programs are strongly NP-hard and NP-hard to approximate within a constant factor. The key link between these two results was proven by Lenstra [16]: mixed integer linear programs can be solved in polynomial time, if the number of integer variables is fixed (but the number of continuous variables does not need to be fixed). In our main result we provide the same type of link between Equations TR and MITR: we prove that an arbitrarily good solution for Equation MITR can be found in polynomial time, if the number $p$ of integer variables is fixed. The key difference between Lenstra’s result and ours, besides the different types of optimization problems considered, is that here we deal with approximate solutions rather than global optimal solutions, but this is unavoidable due to the fact that optimal solutions to Equations TR and MITR are generally irrational. In particular, our findings pave the way for designing efficient trust region methods for mixed integer nonlinear optimization problems.
Ellipsoid-constrained mixed integer quadratic programming. Our main result actually holds for a problem that is more general than Equation MITR. We call it the ellipsoid-constrained mixed integer quadratic programming problem, and we define it as follows:
$\begin{split}\min&\quad x^{\mathsf{T}}Hx+h^{\mathsf{T}}x\\ \operatorname{s.t.}&\quad(x-c)^{\mathsf{T}}Q(x-c)\leq 1\\ &\quad x\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}.\end{split}$ (E-MIQP)
Here $H$ is a symmetric matrix in $\mathbb{Q}^{n\times n}$, $Q$ is a symmetric positive definite matrix in $\mathbb{Q}^{n\times n}$, $h,c\in\mathbb{Q}^{n}$, and $p\in\{0,\dots,n\}$. There are two key differences with Equation MITR: the ball constraint has been replaced with a more general ellipsoid constraint, and the mixed integer lattice $\Pi_{p}(b^{1},\dots,b^{n})+\{c\}$ has been replaced with the “cleaner” mixed integer lattice $\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$.
As we will see in Section 2, Equation MITR is essentially the special case of Equation E-MIQP where the matrix $Q$ is of the form $Q=B^{\mathsf{T}}B$ for some invertible matrix $B\in\mathbb{Q}^{n\times n}$. On the other hand, Equation E-MIQP is indeed more general than Equation MITR, in the sense that there are instances of Equation E-MIQP for which there exists no rational affine transformation that maps them to Equation MITR.
An approximation algorithm for Equation E-MIQP. To state our main results, we first define the concepts of $\epsilon$-approximate solutions and of size. Consider a general optimization problem $\inf\left\{f(x):x\in\mathcal{S}\right\}.$ Let $f_{\inf}$ and $f_{\sup}$ denote the infimum and the supremum of $f(x)$ on $\mathcal{S}$. For $\epsilon\in[0,1]$, we say that $x^{\diamond}\in\mathcal{S}$ is an $\epsilon$-approximate solution if
$f(x^{\diamond})-f_{\inf}\leq\epsilon\cdot(f_{\sup}-f_{\inf}).$
Note that only optimal solutions are 00-approximate solutions, while any point in $\mathcal{S}$ is a $1$-approximate solution. Our definition of approximation has some well-known invariance properties which make it a natural choice for the problems considered in this paper. For instance, it is preserved under dilation and translation of the objective function, and it is insensitive to affine transformations of the objective function and of the feasible region, like for example changes of basis. This definition has been used in our earlier work, and in numerous other places such as [18, 24, 1, 15]. The size (also known as bit size, or length) of rational numbers, vectors, matrices, constraints, and optimization problems, denoted by $\operatorname{size}(\cdot)$, is the standard one in mathematical programming (see, e.g., [22, 4]), and is essentially the number of bits required to encode such objects. We are now ready to state the first main result of this paper.
Theorem 1 (Approximation algorithm for Equation E-MIQP).
There is an algorithm which, for any $\epsilon\in(0,1]$, either finds an $\epsilon$-approximate solution to Equation E-MIQP, or correctly detects that the feasible region is empty. The running time of the algorithm is polynomial in the size of Equation E-MIQP and in $1/\epsilon$, provided that the number $p$ of integer variables is fixed.
Theorem 1 implies that it is significantly simpler to optimize a quadratic function over mixed integer points in an ellipsoid rather than over mixed integer points in a polyhedron, a behavior previously only known in the pure continuous setting [25, 21]. In turn, this suggests that it might be advantageous to design trust region methods for mixed integer nonlinear optimization problems that are based on the 2-norm, or ellipsoidal norms, rather than polyhedral norms.
Even though Equation MITR is a special case of Equation E-MIQP, it still plays a central role in our path to prove Theorem 1. In Sections 4 and 5, we study Equations TR and MITR, and we prove a number of algorithmic and structural results of independent interest. The end goal of these two sections is to prove Proposition 1, which introduces an algorithm for Equation MITR that either finds two feasible solutions whose difference in objective value is at least $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$, for $\epsilon$ that goes to zero, or finds a nonzero vector $d$ along which the feasible region is flat. In Section 6, we get back to the more general Equation E-MIQP. Mapping the ellipsoid defining its feasible region to a new ellipsoid that is arbitrarily close to the unit ball, allows us to employ Proposition 1 to obtain a flatness result for Equation E-MIQP. This result, presented in Proposition 2, gives an algorithm that either finds an $\epsilon$-approximate solution, or finds a nonzero vector $d$ along which the feasible region is flat. In the second case, we are able to map the intersection of the feasible region with a hyperplane, to the feasible region of a new Equation E-MIQP with one fewer integer variable, and as a result we complete our proof of Theorem 1. As we will see, this last step is where the additional level of generality of Equation E-MIQP over Equation MITR plays a key role. For more details on the ideas behind our proofs and techniques used, we refer the reader to the subsequent sections, where some proof overviews are provided to complement the complete and detailed proofs of the main results.
A weak approximation algorithm for Equation MIQP. The techniques that we introduce to prove Theorem 1 are of independent interest and can be used in other mixed integer nonlinear optimization problems. To demonstrate this claim, we consider the mixed integer quadratic programming problem
$\begin{split}\min&\quad x^{\mathsf{T}}Hx+h^{\mathsf{T}}x\\ \operatorname{s.t.}&\quad Wx\leq w\\ &\quad x\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}.\end{split}$ (MIQP)
Here $H$ is a symmetric matrix in $\mathbb{Q}^{n\times n}$, $h\in\mathbb{Q}^{n}$, $W\in\mathbb{Q}^{m\times n}$, $w\in\mathbb{Q}^{m}$, and $p\in\{0,\dots,n\}$. Equation MIQP is NP-hard even in the pure continuous setting [21]. In the general mixed integer case, it is NP-hard to decide whether the feasible region of Equation MIQP is nonempty. However, this feasibility problem can be solved in polynomial time if the number $p$ of integer variables is fixed [16], thus we now focus on this case and assume $p$ is fixed in the subsequent discussion. In [7], the author gives an algorithm that finds an $\epsilon$-approximate solution to Equation MIQP in time polynomial in the size of Equation MIQP and in $1/\epsilon$, provided that the rank of $H$ is fixed. In the general case, where the rank of $H$ is not fixed, no such algorithm can exist, unless P=NP [24]. In fact, unless P=NP, there is a constant $\epsilon\in(0,1)$ such that no algorithm can find an $\epsilon$-approximate solution to Equation MIQP in polynomial time [1]. As a result, we cannot hope to approximate Equation MIQP in polynomial time unless we are willing to accept an approximation factor that tends to one asymptotically as the problem gets larger [25]. In Section 8, we show how the algorithmic and structural results on Equation MITR, and in particular Proposition 1, can be used to design a polynomial time approximation algorithm for Equation MIQP that satisfies a weak approximation bound of this sort.
Theorem 2 (Weak approximation algorithm for Equation MIQP).
Consider Equation MIQP and assume there exists $\bar{f}\in\mathbb{R}$ such that every feasible solution has objective value at least $\bar{f}$. There is an algorithm which either finds an $\epsilon$-approximate solution to the problem, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$, or correctly detects that the feasible region is empty. The running time of the algorithm is polynomial in the size of Equation MIQP, provided that the number $p$ of integer variables is fixed.
Theorem 2 provides an extension to the mixed integer setting of a result by Vavasis in the pure continuous setting [25]. Note that the assumption in Theorem 2 that the objective function is lower bounded on the feasible region cannot be relaxed, unless P=NP [7].
We remark that the main emphasis of this paper lies on the theoretical computational complexity of the algorithms presented. We do not refer to practically efficient implementations of the algorithms, which we believe should be studied in the future. Before proceeding with the proofs of our main results, in Section 2 we discuss the relationship between Equations MITR and E-MIQP, and in Section 3 we study the hardness of the Problems TR, MITR, and E-MIQP.
2 Relationship between Equations MITR and E-MIQP
The next result shows that Equation MITR is essentially the special case of Equation E-MIQP where the matrix $Q$ is of the form $Q=B^{\mathsf{T}}B$ for some invertible matrix $B\in\mathbb{Q}^{n\times n}$.
Observation 1.
Up to a rational affine transformation, Equation MITR is the special case of Equation E-MIQP where the matrix $Q$ is of the form $Q=B^{\mathsf{T}}B$ for some invertible matrix $B\in\mathbb{Q}^{n\times n}$.
Proof.
Consider Equation MITR and perform the change of variables $y:=B^{-1}(x-c)$, where $B$ is the invertible matrix with columns $b^{1},\dots,b^{n}$. The optimization problem that we obtain can be written in the form
$\gamma+\min$ $\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$
$\operatorname{s.t.}$ $\quad(y-c^{\prime})^{\mathsf{T}}Q(y-c^{\prime})\leq 1$
$\quad y\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p},$
where $\gamma:=c^{\mathsf{T}}Hc+h^{\mathsf{T}}c$, $H^{\prime}:=B^{\mathsf{T}}HB$, $h^{\prime}:=2B^{\mathsf{T}}H^{\mathsf{T}}c+B^{\mathsf{T}}h$, $Q:=B^{\mathsf{T}}B$, and $c^{\prime}:=-B^{-1}c$.
Next, consider Equation E-MIQP and assume that the matrix $Q$ is of the form $Q=B^{\mathsf{T}}B$ for some invertible matrix $B\in\mathbb{Q}^{n\times n}$. We perform the change of variables $y:=B(x-c)$ and obtain the optimization problem
$\gamma+\min$ $\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$
$\operatorname{s.t.}$ $\quad y^{\mathsf{T}}y\leq 1$
$\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\},$
where $\gamma:=c^{\mathsf{T}}Hc+{h}^{\mathsf{T}}c$, $H^{\prime}:=B^{-\mathsf{T}}HB^{-1}$, $h^{\prime}:=2B^{-\mathsf{T}}H^{\mathsf{T}}c+B^{-\mathsf{T}}h$, $b^{1},\dots,b^{n}$ are the columns of $B$, and $c^{\prime}:=-Bc$. ∎
Next, we show that Equation E-MIQP is indeed more general than Equation MITR, by presenting instances of Equation E-MIQP for which there exists no rational affine transformation that maps them to Equation MITR. Clearly, there is no rational affine transformation that maps the $1$-dimensional ellipsoid $\left\{x\in\mathbb{R}:2x_{1}^{2}\leq 1\right\}$ to the $1$-dimensional ball $\left\{x\in\mathbb{R}:x_{1}^{2}\leq 1\right\}$. However, this simple example could be written as Equation MITR, if the constraint $x^{\mathsf{T}}x\leq 1$ is replaced with $x^{\mathsf{T}}x\leq r$, for some positive $r\in\mathbb{Q}$. Next, we give a more interesting $2$-dimensional example which shows that Equation E-MIQP is also more general than the extension of Equation MITR with a constraint of the form $x^{\mathsf{T}}x\leq r$, for some positive $r\in\mathbb{Q}$.
Observation 2.
There exists no rational affine transformation that maps the $2$-dimensional ellipsoid $\left\{x\in\mathbb{R}^{2}:2x_{1}^{2}+x_{2}^{2}\leq 1\right\}$ to a $2$-dimensional ball $\left\{x\in\mathbb{R}^{2}:x_{1}^{2}+x_{2}^{2}\leq r\right\}$, for some positive $r\in\mathbb{Q}$.
Proof.
Assume, for a contradiction, that there exists a rational affine transformation as in the statement. The affine transformation is of the form $y=B(x-c)$, for some invertible matrix $B\in\mathbb{Q}^{2\times 2}$, and $c\in\mathbb{Q}^{2}$. We must have $c=0$, since the centers of the ellipsoid and of the ball are both the origin. We then have
$x^{\mathsf{T}}\begin{pmatrix}2r&0\\ 0&r\end{pmatrix}x=y^{\mathsf{T}}y=x^{\mathsf{T}}B^{\mathsf{T}}Bx.$
If we denote by $b_{ij}$ the element of $B$ in position $(i,j)$, these elements must satisfy the system
$\begin{split}&b_{11}^{2}+b_{21}^{2}=2r\\ &b_{12}^{2}+b_{22}^{2}=r\\ &b_{11}b_{12}+b_{21}b_{22}=0.\end{split}$ (1)
From the first two equations in (1), we derive
$2=\frac{b_{11}^{2}+b_{21}^{2}}{b_{12}^{2}+b_{22}^{2}}.$ (2)
From the third equation in (1), we obtain
$b_{11}^{2}b_{12}^{2}=b_{21}^{2}b_{22}^{2}.$
Since $B$ is invertible, $b_{12}$ and $b_{22}$ cannot both equal zero. If $b_{12}\neq 0$, we multiply the right hand of (2) by $b_{12}^{2}/b_{12}^{2}$ and conclude
$2=\frac{b_{11}^{2}b_{12}^{2}+b_{21}^{2}b_{12}^{2}}{b_{12}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{21}^{2}b_{22}^{2}+b_{21}^{2}b_{12}^{2}}{b_{12}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{21}^{2}(b_{12}^{2}+b_{22}^{2})}{b_{12}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{21}^{2}}{b_{12}^{2}}.$
Hence, $\sqrt{2}$ is the rational number $\lvert b_{21}/b_{12}\rvert$, a contradiction. On the other hand, if $b_{22}\neq 0$, we multiply the right hand of (2) by $b_{22}^{2}/b_{22}^{2}$ and conclude
$2=\frac{b_{11}^{2}b_{22}^{2}+b_{21}^{2}b_{22}^{2}}{b_{22}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{11}^{2}b_{22}^{2}+b_{11}^{2}b_{12}^{2}}{b_{22}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{11}^{2}(b_{12}^{2}+b_{22}^{2})}{b_{22}^{2}(b_{12}^{2}+b_{22}^{2})}=\frac{b_{11}^{2}}{b_{22}^{2}}.$
Hence, $\sqrt{2}$ is the rational number $\lvert b_{11}/b_{22}\rvert$, a contradiction. ∎
3 Hardness of mixed integer trust region problems
In this section, we present some hardness results for Problems TR, MITR, and E-MIQP that we mentioned in the introduction. We recall that these three problems are listed in increasing order of generality. We start with a well-known fact about Equation TR.
Observation 3.
There are instances of Equation TR where the optimal value is irrational and no optimal solution is rational, even if $n=2$.
Proof.
Consider the instance of Equation TR with $H:=\begin{pmatrix}2&1\\ 1&0\end{pmatrix}$, and $h:=0$, i.e., $\min\left\{2x_{1}^{2}+2x_{1}x_{2}:x^{\mathsf{T}}x\leq 1\right\}$. It can be checked that the smallest eigenvalue of $H$ is $\lambda_{\min}:=1-\sqrt{2}$. It is well known that $\lambda_{\min}=\min\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x=1\right\}$ and, since $\lambda_{\min}\leq 0$, we also have $\lambda_{\min}=\min\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x\leq 1\right\}$. Hence the optimal value of this instance is $\lambda_{\min}$. This instance has a unique optimal solution, which is the unit length eigenvector $(1-\sqrt{2},1)/\lVert(1-\sqrt{2},1)\rVert_{2}$ associated with the eigenvalue $\lambda_{\min}$. ∎
Observation 3 also holds for the more general Problems MITR, and E-MIQP, even if $p=0$ and $n=2$. These results imply that no polynomial time algorithm, which stores numbers in binary encoding, can find an optimal solution to these problems. Next, we discuss the hardness of Equation MITR. To do so, we need a lemma on the approximation of square roots. While the square root of a rational number can be irrational, it is well known that it can be approximated efficiently via binary search. We formalize this simple result in the next lemma. The proof is given for completeness.
Lemma 1.
Let $\alpha\in\mathbb{Q}$ be positive. There is an algorithm which, for any positive integer $k$, finds nonnegative $l,u\in\mathbb{Q}$ such that $l\leq\sqrt{\alpha}\leq u$ and $u-l=\max\left\{1,\alpha\right\}/2^{k}$. The running time of the algorithm is polynomial in $\operatorname{size}(\alpha),k$.
Proof.
Let $l_{0}:=0$, $u_{0}:=\max\left\{1,\alpha\right\}$. Our algorithm consists of $k$ iterations $i=1,\dots,k$. At iteration $i$, set $m_{i}:=(l_{i-1}+u_{i-1})/2$. Compare $m_{i}^{2}$ with $\alpha$. If $m_{i}^{2}\leq\alpha$, set $l_{i}:=m_{i}$ and $u_{i}:=u_{i-1}$. If $m_{i}^{2}>\alpha$, set $l_{i}:=l_{i-1}$ and $u_{i}:=m_{i}$. When iteration $k$ is completed, return $l:=l_{k}$ and $u:=u_{k}$.
To show correctness of the algorithm, note that $\sqrt{\alpha}\in[0,\max\left\{1,\alpha\right\}]$. The algorithm performs binary search on this interval, comparing, at each iteration, the midpoint $m_{i}$ of the previous interval with $\sqrt{\alpha}$, and updating the interval accordingly. In fact, we have $m_{i}\lesseqgtr\sqrt{\alpha}$ if and only if $m_{i}^{2}\lesseqgtr\alpha$. Hence, for every $i=0,1,\dots,k$, we have $\sqrt{\alpha}\in[l_{i},u_{i}]$. Furthermore, for every $i=0,1,\dots,k$, we have $u_{i}-l_{i}=\max\left\{1,\alpha\right\}/2^{i}$. ∎
Theorem 3.
Equation MITR is strongly NP-hard and it is NP-hard to find an $\epsilon$-approximate solution with $\epsilon<1/17$, even if $p=n$ and the problem is feasible.
Proof.
The Max-Cut problem is strongly NP-hard [13] and it is NP-hard to find an $\epsilon$-approximate solution, with $\epsilon>16/17$ [23]. It is well known that the Max-Cut problem on a graph $G=(V,E)$ with $V=\{1,2,\dots,n\}$ and nonnegative weights $w_{uv}\in\mathbb{Z}$, for $uv\in E$, can be formulated as
$\max$ $\quad\sum_{uv\in E}\frac{w_{uv}}{2}(1-x_{u}x_{v})$
$\operatorname{s.t.}$ $\quad x\in\{-1,1\}^{n}.$
Observe that $\{-1,1\}^{n}=\{x\in 2\mathbb{Z}^{n}+\{1\}:x^{\mathsf{T}}x\leq r\},$ for any $r\in[n,n+8)$. We use Lemma 1 with $\alpha:=n$ and $k:=n$. We obtain $l,u\in\mathbb{Q}$ such that $l\leq\sqrt{n}\leq u$ and $u-l=n/2^{n}$. Therefore, $u-\sqrt{n}\leq n/2^{n}<\sqrt{n+8}-\sqrt{n}$, where the last inequality can be checked with little effort. Hence we have $u\in[\sqrt{n},\sqrt{n+8})$, thus $u^{2}\in[n,n+8)$. Applying the change of variables $y:=x/u$ to the above formulation of the Max-Cut problem, we obtain the following special case of Equation MITR:
$\max$ $\quad\sum_{uv\in E}\frac{w_{uv}}{2}(1-u^{2}y_{u}y_{v})$
$\operatorname{s.t.}$ $\quad y^{\mathsf{T}}y\leq 1$
$\quad y\in\frac{2}{u}\mathbb{Z}^{n}+\left\{\frac{1}{u}\right\}.$
Our next goal is to show that it is NP-hard to decide whether the feasible region of Equation MITR is nonempty. To do so, we first define the concept of lattice. Given linearly independent vectors $b^{1},\dots,b^{p}\in\mathbb{R}^{n}$, we define the lattice
$\Lambda(b^{1},\dots,b^{p}):=\left\{\sum_{i=1}^{p}\mu_{i}b^{i}:\mu_{i}\in\mathbb{Z}\ \forall i=1,\dots,p\right\}.$
Then, we consider the decisional version of the closest vector problem (CVP): Given linearly independent vectors $b^{1},\dots,b^{p}\in\mathbb{Q}^{n}$, a target vector $t\in\mathbb{Q}^{n}$, and $r\in\mathbb{Q}$, determine whether there exists $y\in\Lambda(b^{1},\dots,b^{p})$ with $\lVert y-t\rVert_{2}\leq r$. This problem is well known to be NP-hard (see theorem 3.1 in [17]). We first observe that the problem is NP-hard, even if $p=n$.
Lemma 2.
The decisional version of the CVP is NP-hard, even if $p=n$.
Proof.
We give a polynomial reduction from the decisional version of the CVP. For $j=p+1,\dots,n$, we can find in polynomial time a vector $b^{j}\in\mathbb{Q}^{n}$ orthogonal to $b^{1},\dots,b^{j-1}$ and with $\lVert b^{j}\rVert_{2}>\lVert t\rVert_{2}+r$. This can be done by appropriately rescaling a nonzero solution to the system of linear equations ${b^{i}}^{\mathsf{T}}x=0$, for $i=1,\dots,j-1$. Clearly $b^{1},\dots,b^{n}$ are linearly independent. It is simple to check that the decisional version of the CVP with data $b^{1},\dots,b^{p},t,r$ is equivalent to the decisional version of the CVP with data $b^{1},\dots,b^{n},t,r$. ∎
Theorem 4.
It is NP-hard to decide whether the feasible region of Equation MITR is nonempty, even if $p=n$.
Proof.
We give a polynomial reduction from the decisional version of the CVP with $p=n$. Let $\mathcal{S}:=\{y\in\Lambda(b^{1},\dots,b^{n}):\lVert y-t\rVert_{2}\leq r\}$. Define $x:=(y-t)/r$. We obtain that $\mathcal{S}$ is nonempty if and only if the following set is nonempty:
$\left\{x\in\Lambda\left(\frac{b^{1}}{r},\dots,\frac{b^{n}}{r}\right):\lVert x\rVert_{2}\leq 1\right\}-\left\{\frac{t}{r}\right\}.$
The latter set is the feasible region of Equation MITR with $b^{i}:=b^{i}/r$ for $i=1,\dots,n$, $c:=-t/r$, and $p:=n$. ∎
Since Equation MITR is a special case of Equation E-MIQP, the statements of Theorems 3 and 4 hold also for Equation E-MIQP.
We conclude the section by observing that theorem 1 in [9] implies that the problem of deciding whether the feasible region of Equation E-MIQP is nonempty lies in the complexity class NP. Since Equation MITR can be mapped to Equation E-MIQP with the change of variables $y:=B^{-1}(x-c)$, where $B$ is the invertible matrix with columns $b^{1},\dots,b^{n}$, we obtain that also the problem of deciding whether the feasible region of Equation MITR is nonempty lies in the complexity class NP. Theorem 4 then implies that these two decision problems are NP-complete.
4 Algorithmic and structural results for Equation TR
In this section we obtain a number of algorithmic and structural results for Equation TR which we need to prove our main theorems. In order to streamline the presentation, we introduce some notation. A ball with center $c$ and radius $r$ is a set of the form
$\mathcal{B}(c,r):=\{x\in\mathbb{R}^{n}:\lVert x-c\rVert_{2}\leq r\},$
where $c\in\mathbb{R}^{n}$, and $r\in\mathbb{R}$ is positive. Notice that the feasible region of Equation TR is the ball $\mathcal{B}(0,1)$. For ease of notation, from now on, we denote the objective function of Equation TR by
$f(x):=x^{\mathsf{T}}Hx+h^{\mathsf{T}}x.$
In this paper, we denote by $\log(x)$ the logarithm to the base $2$ of a positive $x\in\mathbb{R}$.
4.1 Algorithmic results
Vavasis and Zippel [26] showed that we can find an arbitrarily good solution to Equation TR in polynomial time. In this section, we discuss some algorithmic consequences of this result.
Lemma 3 ([26]).
Let $x^{\circ}$ be an optimal solution of Equation TR. There is an algorithm which, for any positive integer $k$, finds a feasible solution $x^{\dagger}$ to Equation TR such that
$f(x^{\dagger})-f(x^{\circ})\leq 2^{-k}.$
The running time of the algorithm is polynomial in the size of Equation TR and in $k$.
Proof.
Let $x^{\dagger}$ be the feasible solution to Equation TR obtained after $k^{\prime}$ iterations with the algorithm in [27, 26]. From [26] (see in particular the first three lines of section 5), we know $f(x^{\dagger})-f(x^{\circ})\leq 2^{q-k^{\prime}}$, where $q$ is a polynomial in $\operatorname{size}(H)$ and $\operatorname{size}(h)$, whose explicit form can be derived from section 4 in [26]. The result follows by setting $k^{\prime}:=\lceil k+q\rceil$, since each iteration of the algorithm performs a number of operations that is polynomial in $\operatorname{size}(H)$ and $\operatorname{size}(h)$. ∎
Next, we restate Lemma 3 in such a way that the upper bound on the gap $f(x^{\dagger})-f(x^{\circ})$ is of the form $\epsilon\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$. Given a matrix $A\in\mathbb{R}^{m\times n}$, we denote by $\lVert A\rVert_{2}$ its spectral norm, which is defined by $\lVert A\rVert_{2}:=\sup_{x\neq 0}\lVert Ax\rVert_{2}/\lVert x\rVert_{2}$. If $A$ is symmetric, $\lVert A\rVert_{2}$ is equal to the maximum absolute value among the eigenvalues of $A$.
Lemma 4.
Let $x^{\circ}$ be an optimal solution of Equation TR. There is an algorithm which, for any positive $\epsilon\in\mathbb{Q}$, finds a feasible solution $x^{\dagger}$ to Equation TR such that
$f(x^{\dagger})-f(x^{\circ})\leq\epsilon\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}.$
The running time of the algorithm is polynomial in the size of Equation TR and in $\log(1/\epsilon)$.
Proof.
If $H$ is the zero matrix and $h$ is the zero vector, we can simply set $x^{\dagger}:=0$. Thus, in the remainder of the proof we assume that at least one among $H$ and $h$ is nonzero. We set $k:=\lceil\log(1/\epsilon)+\operatorname{size}(H)+\operatorname{size}(h)\rceil$, which can be computed in polynomial time. From Lemma 3, it suffices to show
$2^{-k}\leq\epsilon\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}.$
We consider separately two cases. In the first case, we assume $\lVert h\rVert_{2}\geq\lVert H\rVert_{2}$. Then $h$ is nonzero, thus at least one component is nonzero, say $h_{i}$. We obtain
$\lVert h\rVert_{2}=\sqrt{h_{1}^{2}+\cdots+h_{n}^{2}}\geq\lvert h_{i}\rvert\geq 2^{-\operatorname{size}(h)}.$
Hence,
$2^{-k}\leq 2^{-\log(1/\epsilon)-\operatorname{size}(h)}=2^{-\log(1/\epsilon)}2^{-\operatorname{size}(h)}\leq\epsilon\lVert h\rVert_{2}=\epsilon\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}.$
In the second case, we assume $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$. Then $H$ is nonzero, thus at least one element is nonzero, say $h_{ij}$. Let $e^{j}$ be the $j$th vector of the standard basis of $\mathbb{R}^{n}$. We obtain
$\lVert H\rVert_{2}=\sup_{x\neq 0}\frac{\lVert Hx\rVert_{2}}{\lVert x\rVert_{2}}\geq\frac{\lVert He^{j}\rVert_{2}}{\lVert e^{j}\rVert_{2}}=\sqrt{h_{1j}^{2}+\cdots+h_{nj}^{2}}\geq\lvert h_{ij}\rvert\geq 2^{-\operatorname{size}(H)}.$
Hence,
$2^{-k}\leq 2^{-\log(1/\epsilon)-\operatorname{size}(H)}=2^{-\log(1/\epsilon)}2^{-\operatorname{size}(H)}\leq\epsilon\lVert H\rVert_{2}=\epsilon\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}.$
We now show how Lemma 4 can be used to compute a vector $v\in\mathcal{B}(0,1)$ with $|v^{\mathsf{T}}Hv|$ arbitrarily close to the spectral norm of $H$.
Lemma 5.
Let $H$ be a symmetric matrix in $\mathbb{Q}^{n\times n}$. There is an algorithm which, for any positive $\epsilon\in\mathbb{Q}$, finds a vector $v\in\mathcal{B}(0,1)$ such that
$(1-\epsilon)\lVert H\rVert_{2}\leq|v^{\mathsf{T}}Hv|\leq\lVert H\rVert_{2}.$
The running time of the algorithm is polynomial in $\operatorname{size}(H)$, $\log(1/\epsilon)$.
Proof.
If $H$ is the zero matrix, we can simply set $v:=0$, thus in the remainder of the proof we assume $H$ nonzero. Consider the following two optimization problems, which are special cases of Equation TR:
$(S):=\ \min$ $\quad x^{\mathsf{T}}Hx$ $(T):=\ \max$ $\quad x^{\mathsf{T}}Hx$ $=\ -\min$ $\quad x^{\mathsf{T}}(-H)x$
$\operatorname{s.t.}$ $\quad x^{\mathsf{T}}x\leq 1,$ $\operatorname{s.t.}$ $\quad x^{\mathsf{T}}x\leq 1$ $\operatorname{s.t.}$ $\quad x^{\mathsf{T}}x\leq 1.$
The algorithm is defined as follows. We apply Lemma 4 to both Problems $(S)$ and $(T)$, and find a feasible solution $s$ for $(S)$ and a feasible solution $t$ for $(T)$. Return $v:=t$ if $\lvert{t}^{\mathsf{T}}H{t}\rvert\geq\lvert{s}^{\mathsf{T}}H{s}\rvert$, and return $v:={s}$ otherwise. In the remainder of the proof, we show the correctness of the algorithm.
Denote by $\lambda_{\min}$ the smallest eigenvalue of $H$ and by $\lambda_{\max}$ the largest eigenvalue of $H$. Since $H$ is symmetric, we have $\lambda_{\min}=\min\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x=1\right\}$ and $\lambda_{\max}=\max\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x=1\right\}$. Since $H$ is nonzero, we have $\lVert H\rVert_{2}=\max\{|\lambda_{\min}|,\ |\lambda_{\max}|\}>0$. Every $x\in\mathbb{R}^{n}$ with $x^{\mathsf{T}}x\leq 1$ satisfies $|x^{\mathsf{T}}Hx|\leq\lVert H\rVert_{2}$, thus in the remainder of the proof it suffices to show $|v^{\mathsf{T}}Hv|\geq(1-\epsilon)\lVert H\rVert_{2}$. From the definition of our algorithm, it suffices to show either $|{s}^{\mathsf{T}}H{s}|\geq(1-\epsilon)\lVert H\rVert_{2}$ or $|{t}^{\mathsf{T}}H{t}|\geq(1-\epsilon)\lVert H\rVert_{2}$. In the rest of the proof, we consider separately two cases.
In the first case, we assume $\lVert H\rVert_{2}=|\lambda_{\min}|$. We then have $\lambda_{\min}<0$. Hence, $\lambda_{\min}=\min\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x\leq 1\right\}$. From Lemma 4, we know ${s}^{\mathsf{T}}Hs-\lambda_{\min}\leq\epsilon\lVert H\rVert_{2}$. As $\lambda_{\min}=-\lVert H\rVert_{2}$, this gives $|{s}^{\mathsf{T}}H{s}|\geq-{s}^{\mathsf{T}}H{s}\geq(1-\epsilon)\lVert H\rVert_{2}$, as desired.
Next, we consider the case $\lVert H\rVert_{2}=|\lambda_{\max}|$. We then have $\lambda_{\max}>0$. Hence, $\lambda_{\max}=\max\left\{x^{\mathsf{T}}Hx:x^{\mathsf{T}}x\leq 1\right\}$. From Lemma 4, we know $-{t}^{\mathsf{T}}Ht+\lambda_{\max}\leq\epsilon\lVert H\rVert_{2}$. As $\lambda_{\max}=\lVert H\rVert_{2}$, this gives $|{t}^{\mathsf{T}}H{t}|\geq{t}^{\mathsf{T}}H{t}\geq(1-\epsilon)\lVert H\rVert_{2}$, as desired. ∎
4.2 Feasible points to Equation TR with distant objective values
In this section, we provide two lower bounds on the difference, in objective value, between some specific feasible points to Equation TR. First, we show that two vectors $s,t$, arbitrarily close to $h/\lVert h\rVert_{2}$ and to $-h/\lVert h\rVert_{2}$, respectively, yield a lower bound on $f(s)-f(t)$ that tends to $2\lVert h\rVert_{2}$, as $\epsilon$ goes to zero. We remark that the bounds that we present are of interest when $\epsilon$ is close to zero.
Lemma 6.
Let $H$ be a symmetric matrix in $\mathbb{R}^{n\times n}$, let $h\in\mathbb{R}^{n}$, and let $\epsilon\in[0,1/2]$. Assume $\lVert H\rVert_{2}\leq\lVert h\rVert_{2}$. Let $l\in\mathbb{R}$ be positive such that $1-\epsilon\leq l\lVert h\rVert_{2}\leq 1$. Let $s\in\mathcal{B}((1-\epsilon)lh,\epsilon)$ and $t\in\mathcal{B}(-(1-\epsilon)lh,\epsilon)$. Then
$f(s)-f(t)\geq 2\left(1-5\epsilon\right)\lVert h\rVert_{2}.$
Proof.
To show the result, we derive a lower bound for $f(s)$ and an upper bound for $f(t)$. First, we give a lower bound for $f(s)$. Let $g\in\mathbb{R}^{n}$ so that $s=(1-\epsilon)lh+g$. Since $s\in\mathcal{B}((1-\epsilon)lh,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
$f(s)$ $={s}^{\mathsf{T}}Hs+h^{\mathsf{T}}s$
$=((1-\epsilon)lh+g)^{\mathsf{T}}H((1-\epsilon)lh+g)+h^{\mathsf{T}}((1-\epsilon)lh+g)$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+g^{\mathsf{T}}Hg+2(1-\epsilon)lh^{\mathsf{T}}Hg+(1-\epsilon)l\lVert h\rVert_{2}^{2}+h^{\mathsf{T}}g$
$\geq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\lvert g^{\mathsf{T}}Hg\rvert-\lvert 2(1-\epsilon)lh^{\mathsf{T}}Hg\rvert+(1-\epsilon)l\lVert h\rVert_{2}^{2}-\lvert h^{\mathsf{T}}g\rvert$
$\geq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}-2(1-\epsilon)l\lVert h\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}+(1-\epsilon)l\lVert h\rVert_{2}^{2}-\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\geq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\epsilon^{2}\lVert h\rVert_{2}-2\epsilon(1-\epsilon)l\lVert h\rVert_{2}^{2}+(1-\epsilon)l\lVert h\rVert_{2}^{2}-\epsilon\lVert h\rVert_{2}$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\epsilon^{2}\lVert h\rVert_{2}+(1-2\epsilon)(1-\epsilon)l\lVert h\rVert_{2}^{2}-\epsilon\lVert h\rVert_{2}$
$\geq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\epsilon^{2}\lVert h\rVert_{2}+(1-2\epsilon)(1-\epsilon)^{2}\lVert h\rVert_{2}-\epsilon\lVert h\rVert_{2}$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\left(1-5\epsilon+4\epsilon^{2}-2\epsilon^{3}\right)\lVert h\rVert_{2}.$
Here, in the third inequality we used $\lVert g\rVert_{2}\leq\epsilon$, $\lVert H\rVert_{2}\leq\lVert h\rVert_{2}$, and in the fourth inequality $l\lVert h\rVert_{2}\geq 1-\epsilon$, since $1-2\epsilon\geq 0$.
Similarly, we give an upper bound for $f(t)$. Redefine $g\in\mathbb{R}^{n}$ so that $t=-(1-\epsilon)lh+g$. Since $t\in\mathcal{B}(-(1-\epsilon)lh,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
$f(t)$ $={t}^{\mathsf{T}}Ht+h^{\mathsf{T}}t$
$=(-(1-\epsilon)lh+g)^{\mathsf{T}}H(-(1-\epsilon)lh+g)+h^{\mathsf{T}}(-(1-\epsilon)lh+g)$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+g^{\mathsf{T}}Hg-2(1-\epsilon)lh^{\mathsf{T}}Hg-(1-\epsilon)l\lVert h\rVert_{2}^{2}+h^{\mathsf{T}}g$
$\leq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\lvert g^{\mathsf{T}}Hg\rvert+\lvert 2(1-\epsilon)lh^{\mathsf{T}}Hg\rvert-(1-\epsilon)l\lVert h\rVert_{2}^{2}+\lvert h^{\mathsf{T}}g\rvert$
$\leq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}+2(1-\epsilon)l\lVert h\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}-(1-\epsilon)l\lVert h\rVert_{2}^{2}+\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\leq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\epsilon^{2}\lVert h\rVert_{2}+2\epsilon(1-\epsilon)l\lVert h\rVert_{2}^{2}-(1-\epsilon)l\lVert h\rVert_{2}^{2}+\epsilon\lVert h\rVert_{2}$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\epsilon^{2}\lVert h\rVert_{2}-(1-2\epsilon)(1-\epsilon)l\lVert h\rVert_{2}^{2}+\epsilon\lVert h\rVert_{2}$
$\leq(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh+\epsilon^{2}\lVert h\rVert_{2}-(1-2\epsilon)(1-\epsilon)^{2}\lVert h\rVert_{2}+\epsilon\lVert h\rVert_{2}$
$=(1-\epsilon)^{2}l^{2}h^{\mathsf{T}}Hh-\left(1-5\epsilon+4\epsilon^{2}-2\epsilon^{3}\right)\lVert h\rVert_{2}.$
Here, in the third inequality we used $\lVert g\rVert_{2}\leq\epsilon$, $\lVert H\rVert_{2}\leq\lVert h\rVert_{2}$, and in the fourth inequality $l\lVert h\rVert_{2}\geq 1-\epsilon$, since $1-2\epsilon\geq 0$.
We combine the two bounds above and obtain
$f(s)-f(t)$ $\geq 2\left(1-5\epsilon+4\epsilon^{2}-2\epsilon^{3}\right)\lVert h\rVert_{2}$
$\geq 2\left(1-5\epsilon\right)\lVert h\rVert_{2}.$
Next, we show that two vectors $u,w$, arbitrarily close to the origin, and to a vector $v$ as in Lemma 5, yield a lower bound on $\lvert f(w)-f(u)\rvert$ that tends to $\lVert H\rVert_{2}$, as $\epsilon$ goes to zero.
Lemma 7.
Let $H$ be a symmetric matrix in $\mathbb{R}^{n\times n}$, let $h\in\mathbb{R}^{n}$, and let $\epsilon\in[0,1]$. Assume $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. Let $v\in\mathcal{B}(0,1)$ such that $|v^{\mathsf{T}}Hv|\geq(1-\epsilon)\lVert H\rVert_{2}$. Assume $h^{\mathsf{T}}v\geq 0$ if $v^{\mathsf{T}}Hv\geq 0$, and $h^{\mathsf{T}}v\leq 0$ if $v^{\mathsf{T}}Hv<0$. Let $w\in\mathcal{B}((1-\epsilon)v,\epsilon)$ and $u\in\mathcal{B}(0,\epsilon)$. Then
$\lvert f(w)-f(u)\rvert\geq\left(1-7\epsilon\right)\lVert H\rVert_{2}.$
Proof.
First, we consider the case $v^{\mathsf{T}}Hv\geq 0$. To show the result, we derive an upper bound for $f(u)$ and a lower bound for $f(w)$. We start with an upper bound for $f(u)$.
$f(u)$ $={u}^{\mathsf{T}}Hu+h^{\mathsf{T}}u$
$\leq\lvert{u}^{\mathsf{T}}Hu\rvert+\lvert h^{\mathsf{T}}u\rvert$
$\leq\lVert u\rVert_{2}^{2}\lVert H\rVert_{2}+\lVert h\rVert_{2}\lVert u\rVert_{2}$
$\leq\epsilon^{2}\lVert H\rVert_{2}+\epsilon\lVert H\rVert_{2}$
$=\left(\epsilon+\epsilon^{2}\right)\lVert H\rVert_{2}.$
Here, the third inequality holds because $\lVert u\rVert_{2}\leq\epsilon$ and $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. Next, we give a lower bound for $f(w)$. Let $g\in\mathbb{R}^{n}$ so that $w=(1-\epsilon)v+g$. Since $w\in\mathcal{B}((1-\epsilon)v,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
$f(w)$ $={w}^{\mathsf{T}}Hw+h^{\mathsf{T}}w$
$=((1-\epsilon)v+g)^{\mathsf{T}}H((1-\epsilon)v+g)+h^{\mathsf{T}}((1-\epsilon)v+g)$
$=(1-\epsilon)^{2}v^{\mathsf{T}}Hv+2(1-\epsilon)v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+(1-\epsilon)h^{\mathsf{T}}v+h^{\mathsf{T}}g$
$\geq(1-\epsilon)^{2}v^{\mathsf{T}}Hv+2(1-\epsilon)v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+h^{\mathsf{T}}g$
$\geq(1-\epsilon)^{2}v^{\mathsf{T}}Hv-\lvert 2(1-\epsilon)v^{\mathsf{T}}Hg\rvert-\lvert g^{\mathsf{T}}Hg\rvert-\lvert h^{\mathsf{T}}g\rvert$
$\geq(1-\epsilon)^{3}\lVert H\rVert_{2}-2(1-\epsilon)\lVert v\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}-\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}-\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\geq(1-\epsilon)^{3}\lVert H\rVert_{2}-2\epsilon(1-\epsilon)\lVert H\rVert_{2}-\epsilon^{2}\lVert H\rVert_{2}-\epsilon\lVert H\rVert_{2}$
$=\left(1-6\epsilon+4\epsilon^{2}-\epsilon^{3}\right)\lVert H\rVert_{2}.$
Here, in the first inequality we used the assumption $h^{\mathsf{T}}v\geq 0$, in the third inequality $v^{\mathsf{T}}Hv\geq(1-\epsilon)\lVert H\rVert_{2}$, and in the fourth inequality $\lVert v\rVert_{2}\leq 1$, $\lVert g\rVert_{2}\leq\epsilon$, $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. We combine the two bounds above and obtain
$f(w)-f(u)$ $\geq\left(1-7\epsilon+3\epsilon^{2}-\epsilon^{3}\right)\lVert H\rVert_{2}$
$\geq\left(1-7\epsilon\right)\lVert H\rVert_{2}.$
Next, we consider the case $v^{\mathsf{T}}Hv<0$. This case is very similar to the previous one, and we provide a proof for completeness. To show the result, we derive a lower bound for $f(u)$ and an upper bound for $f(w)$. We start with a lower bound for $f(u)$.
$f(u)$ $={u}^{\mathsf{T}}Hu+h^{\mathsf{T}}u$
$\geq-\lvert{u}^{\mathsf{T}}Hu\rvert-\lvert h^{\mathsf{T}}u\rvert$
$\geq-\lVert u\rVert_{2}^{2}\lVert H\rVert_{2}-\lVert h\rVert_{2}\lVert u\rVert_{2}$
$\geq-\epsilon^{2}\lVert H\rVert_{2}-\epsilon\lVert H\rVert_{2}$
$=-\left(\epsilon+\epsilon^{2}\right)\lVert H\rVert_{2}.$
Here, the third inequality holds because $\lVert u\rVert_{2}\leq\epsilon$ and $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. Next, we give an upper bound for $f(w)$. Let $g\in\mathbb{R}^{n}$ so that $w=(1-\epsilon)v+g$. Since $w\in\mathcal{B}((1-\epsilon)v,\epsilon)$, we have $\lVert g\rVert_{2}\leq\epsilon$.
$f(w)$ $={w}^{\mathsf{T}}Hw+h^{\mathsf{T}}w$
$=((1-\epsilon)v+g)^{\mathsf{T}}H((1-\epsilon)v+g)+h^{\mathsf{T}}((1-\epsilon)v+g)$
$=(1-\epsilon)^{2}v^{\mathsf{T}}Hv+2(1-\epsilon)v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+(1-\epsilon)h^{\mathsf{T}}v+h^{\mathsf{T}}g$
$\leq(1-\epsilon)^{2}v^{\mathsf{T}}Hv+2(1-\epsilon)v^{\mathsf{T}}Hg+g^{\mathsf{T}}Hg+h^{\mathsf{T}}g$
$\leq(1-\epsilon)^{2}v^{\mathsf{T}}Hv+\lvert 2(1-\epsilon)v^{\mathsf{T}}Hg\rvert+\lvert g^{\mathsf{T}}Hg\rvert+\lvert h^{\mathsf{T}}g\rvert$
$\leq-(1-\epsilon)^{3}\lVert H\rVert_{2}+2(1-\epsilon)\lVert v\rVert_{2}\lVert H\rVert_{2}\lVert g\rVert_{2}+\lVert g\rVert_{2}^{2}\lVert H\rVert_{2}+\lVert h\rVert_{2}\lVert g\rVert_{2}$
$\leq-(1-\epsilon)^{3}\lVert H\rVert_{2}+2\epsilon(1-\epsilon)\lVert H\rVert_{2}+\epsilon^{2}\lVert H\rVert_{2}+\epsilon\lVert H\rVert_{2}$
$=-\left(1-6\epsilon+4\epsilon^{2}-\epsilon^{3}\right)\lVert H\rVert_{2}.$
Here, in the first inequality we used the assumption $h^{\mathsf{T}}v\leq 0$, in the third inequality $v^{\mathsf{T}}Hv\leq-(1-\epsilon)\lVert H\rVert_{2}$, and in the fourth inequality $\lVert v\rVert_{2}\leq 1$, $\lVert g\rVert_{2}\leq\epsilon$, $\lVert h\rVert_{2}\leq\lVert H\rVert_{2}$. We combine the two bounds above and obtain
$f(u)-f(w)$ $\geq\left(1-7\epsilon+3\epsilon^{2}-\epsilon^{3}\right)\lVert H\rVert_{2}$
$\geq\left(1-7\epsilon\right)\lVert H\rVert_{2}.$
5 Feasible points to Equation MITR with distant objective values
In this section we study Equation MITR. The main goal is to introduce an algorithm that either finds two feasible solutions whose difference in objective value is at least $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$, for $\epsilon$ that goes to zero, or finds a nonzero vector $d$ along which the feasible region is flat. This algorithm will be a key component in the design of the approximation algorithms for Equations E-MIQP and MIQP stated in Theorems 1 and 2. The reason we seek a bound featuring $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$ will become clear later, in the proof of Proposition 2, and is essentially because the same $\max$ appears in the bound in Lemma 4. To state the main result of this section, we need to introduce the concept of width. Let $\mathcal{S}\subseteq\mathbb{R}^{n}$ be a nonempty bounded closed set and let $d\in\mathbb{R}^{n}$. We define the width of $\mathcal{S}$ along $d$ to be
$\operatorname{width}_{d}(\mathcal{S}):=\max\left\{d^{\mathsf{T}}x:x\in\mathcal{S}\right\}-\min\left\{d^{\mathsf{T}}x:x\in\mathcal{S}\right\}.$
Proposition 1 (Distant points for Equation MITR).
There is an algorithm which, for any $\epsilon\in(0,1/3]$, either finds two feasible solutions $x^{1},x^{2}$ to Equation MITR such that
$f(x^{1})-f(x^{2})\geq\left(1-7\epsilon\right)\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\},$
or finds a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}(0,1))\leq p2^{p(p-1)/4}/\epsilon$. The running time of the algorithm is polynomial in the size of Equation MITR and in $\log(1/\epsilon)$.
5.1 Ingredients
In order to prove Proposition 1, we need two more ingredients. Our first ingredient concerns the approximation of square roots, and is a consequence of Lemma 1. The key difference is that the approximation gap is multiplicative, rather than additive.
Lemma 8.
Let $\alpha\in\mathbb{Q}$ be positive. There is an algorithm which, for any positive $\epsilon\in\mathbb{Q}$, finds a positive $l\in\mathbb{Q}$ such that $l\leq\sqrt{\alpha}\leq(1+\epsilon)l$. The running time of the algorithm is polynomial in $\operatorname{size}(\alpha),\log(1/\epsilon)$.
Proof.
We use Lemma 1 with $k:=\lceil\log(1/\epsilon)+3\operatorname{size}(\alpha)/2+1\rceil$. We only need to show that the obtained $l,u$ satisfy $l>0$ and $u\leq(1+\epsilon)l$.
Using $k\geq 3\operatorname{size}(\alpha)/2+1$ and $2^{-\operatorname{size}(\alpha)}\leq\alpha\leq 2^{\operatorname{size}(\alpha)}$, we obtain
$\begin{split}l&=u-\frac{\max\left\{1,\alpha\right\}}{2^{k}}\geq\sqrt{\alpha}-\frac{\max\left\{1,\alpha\right\}}{2^{3\operatorname{size}(\alpha)/2+1}}\geq 2^{-\operatorname{size}(\alpha)/2}-\frac{2^{\operatorname{size}(\alpha)}}{2^{3\operatorname{size}(\alpha)/2+1}}\\ &=2^{-\operatorname{size}(\alpha)/2}-2^{-\operatorname{size}(\alpha)/2-1}=2^{-\operatorname{size}(\alpha)/2-1},\end{split}$ (3)
which implies $l>0$.
We now show $u\leq(1+\epsilon)l$. Using $k\geq\log(1/\epsilon)+3\operatorname{size}(\alpha)/2+1$, $\alpha\leq 2^{\operatorname{size}(\alpha)}$, and (3) we have
$u$ $=l+\frac{\max\left\{1,\alpha\right\}}{2^{k}}\leq l+\frac{2^{\operatorname{size}(\alpha)}}{2^{\log(1/\epsilon)+3\operatorname{size}(\alpha)/2+1}}$
$=l+2^{-\operatorname{size}(\alpha)/2-1}2^{\log(\epsilon)}\leq l+l\epsilon=(1+\epsilon)l.$
The second ingredient is a flatness result for balls, which follows with little effort from proposition 4 in [7].
Lemma 9 (Flatness lemma).
Let $a,c\in\mathbb{Q}^{n}$, let $\delta\in\mathbb{Q}$ with $\delta\geq 0$, let $b^{1},\dots,b^{n}$ be linearly independent vectors in $\mathbb{Q}^{n}$, and let $p\in\{0,\dots,n\}$. There is a polynomial time algorithm which either finds a vector in $\mathcal{B}(a,\delta)\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or finds a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}(a,\delta))\leq p2^{p(p-1)/4}$.
Proof.
In this proof, we denote by $\operatorname{span}(\mathcal{S})$ the linear span of a set $\mathcal{S}$ of vectors and by $\mathcal{L}^{\perp}$ the orthogonal complement of a linear subspace $\mathcal{L}$ or $\mathbb{R}^{n}$. For $i=1,\dots,p$, let $c^{i}$ be the orthogonal projection of $b^{i}$ on $\operatorname{span}(b^{p+1},\dots,b^{n})^{\perp}$, i.e., $c^{i}:=b^{i}-\sum_{j=p+1}^{n}({b^{j}}^{\mathsf{T}}b^{i}/\lVert b^{j}\rVert_{2}^{2})b^{j}$. We apply proposition 4 in [7] with $\Lambda:=\Lambda(c^{1},\dots,c^{p})$, $d:=n$, and $a:=a-c$. Then, there is a polynomial time algorithm which either finds a vector $\bar{x}\in\mathcal{B}(a-c,\delta)\cap(\Lambda+\operatorname{span}(\Lambda)^{\perp})$, or finds a nonzero vector $d\in\operatorname{span}(\Lambda)$ with $d^{\mathsf{T}}c^{1},\dots,d^{\mathsf{T}}c^{p}$ integer such that $\operatorname{width}_{d}(\mathcal{B}(a-c,\delta))\leq p2^{p(p-1)/4}$.
In the first case, the vector $\bar{x}^{\prime}:=\bar{x}+c$ is in $\mathcal{B}(a,\delta)\cap(\Lambda+\operatorname{span}(\Lambda)^{\perp}+\{c\})$, and we have $\Lambda+\operatorname{span}(\Lambda)^{\perp}=\Lambda+\operatorname{span}(b^{p+1},\dots,b^{n})=\Pi_{p}(b^{1},\dots,b^{n})$. In the second case, $d\in\operatorname{span}(\Lambda)$ implies $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$. We then have $d^{\mathsf{T}}c^{i}=d^{\mathsf{T}}b^{i}$, for $i=1,\dots,p$, thus $d^{\mathsf{T}}c^{1},\dots,d^{\mathsf{T}}c^{p}$ integer implies $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer. Furthermore, by definition of width, we obtain $\operatorname{width}_{d}(\mathcal{B}(a,\delta))=\operatorname{width}_{d}(\mathcal{B}(a-c,\delta))\leq p2^{p(p-1)/4}$. ∎
5.2 Proof of Proposition 1
We are now ready to show Proposition 1. Before presenting its proof, we briefly discuss the the idea of the algorithm. The goal is to either find two feasible solutions to Equation MITR whose difference in objective value is at least $\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}$, for $\epsilon$ that goes to zero, or to find a nonzero vector $d$ along which the feasible region is flat. Depending on which one among $\lVert H\rVert_{2}$ and $\lVert h\rVert_{2}$ is largest, we search for the two feasible solutions in different regions in $\mathcal{B}(0,1)$; As a result, we consider separately the cases $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$ and $\lVert h\rVert_{2}>\lVert H\rVert_{2}$.
In the case $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$, we would like to consider a vector $v\in\mathcal{B}(0,1)$ with $|v^{\mathsf{T}}Hv|=\lVert H\rVert_{2}$. However, this vector could be irrational, thus we use Lemma 5 to find, instead, a vector $v\in\mathcal{B}(0,1)$ with $|v^{\mathsf{T}}Hv|$ that is $\epsilon$-close to $\lVert H\rVert_{2}$. Next, we use our flatness lemma to seek two feasible points $u,w$ to Equation MITR that are $\epsilon$-close to the origin and to $v$, respectively. If we find both these vectors, we use Lemma 7 to show that they have a difference in objective value that is at least $\lVert H\rVert_{2}$, for $\epsilon$ that goes to zero. Otherwise, we show that we have found a nonzero vector ($d$ in the statement) along which the feasible region $\mathcal{B}(0,1)$ is flat.
In the remaining case $\lVert h\rVert_{2}>\lVert H\rVert_{2}$, we would like to consider the vector $h/\lVert h\rVert_{2}$. However, $\lVert h\rVert_{2}$ could be irrational, thus we find instead a scalar $l$ such that $lh\in\mathcal{B}(0,1)$ is $\epsilon$-close to $h/\lVert h\rVert_{2}$. Next, we use our flatness lemma to seek two feasible points $s,t$ to Equation MITR that are $\epsilon$-close to $lh$ and to $-lh$, respectively. If we find both these vectors, we use Lemma 6 to show that they have a difference in objective value that is at least $\lVert h\rVert_{2}$, for $\epsilon$ that goes to zero. Otherwise, we show that we have found a nonzero vector ($d$ in the statement) along which the feasible region $\mathcal{B}(0,1)$ is flat. We now give the complete proof of Proposition 1, which uses Lemmas 8, 5, 7, 6 and 9.
Proof.
In this proof, we consider separately the cases $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$ and $\lVert h\rVert_{2}>\lVert H\rVert_{2}$.
We start with the case $\lVert H\rVert_{2}\geq\lVert h\rVert_{2}$. Let $\mathcal{B}_{u}:=\mathcal{B}(0,\epsilon)$. From Lemma 9 (with $a:=0$ and $\delta:=\epsilon$), we either find a vector $u\in\mathcal{B}_{u}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{u})\leq p2^{p(p-1)/4}$. Using the algorithm in Lemma 5, we find a vector $v\in\mathcal{B}(0,1)$ such that $|v^{\mathsf{T}}Hv|\geq(1-\epsilon)\lVert H\rVert_{2}$. Note that the vector $-v$ also lies in $\mathcal{B}(0,1)$ and we have $(-v)^{\mathsf{T}}H(-v)=v^{\mathsf{T}}Hv$. Hence, by possibly redefining $v:=-v$, we now assume $h^{\mathsf{T}}v\geq 0$ if $v^{\mathsf{T}}Hv\geq 0$, and $h^{\mathsf{T}}v\leq 0$ if $v^{\mathsf{T}}Hv<0$. Let $\mathcal{B}_{w}:=\mathcal{B}((1-\epsilon)v,\epsilon)$. From Lemma 9, we either find a vector $w\in\mathcal{B}_{w}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{w})\leq p2^{p(p-1)/4}$.
Consider the subcase where Lemma 9 successfully found both vectors $u$ and $w$. Note that $\mathcal{B}_{u}$ and $\mathcal{B}_{w}$ are contained in $\mathcal{B}(0,1)$, thus $u,w$ are both feasible to Equation MITR. We apply Lemma 7 and obtain
$\lvert f(w)-f(u)\rvert\geq\left(1-7\epsilon\right)\lVert H\rVert_{2}.$
Next, consider the subcase where Lemma 9 did not find both vectors $u$ and $w$. In this case, Lemma 9 found a vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B})\leq p2^{p(p-1)/4}$, for $\mathcal{B}\in\{\mathcal{B}_{u},\mathcal{B}_{w}\}$. We then have
$\operatorname{width}_{d}(\mathcal{B}(0,1))$ $=\frac{1}{\epsilon}\operatorname{width}_{d}(\mathcal{B})\leq\frac{1}{\epsilon}p2^{p(p-1)/4}.$
This concludes the proof in the first case.
In the remainder of the proof, we consider the case $\lVert h\rVert_{2}>\lVert H\rVert_{2}$. In particular, we have $h\neq 0$. Using the algorithm in Lemma 8 (with $\alpha:=1/\lVert h\rVert_{2}^{2}$), we find a positive $l\in\mathbb{Q}$ such that $l\leq 1/\lVert h\rVert_{2}\leq(1+\epsilon)l$. We obtain $l\lVert h\rVert_{2}\leq 1$ and $l\lVert h\rVert_{2}\geq 1/(1+\epsilon)\geq 1-\epsilon$, thus $1-\epsilon\leq l\lVert h\rVert_{2}\leq 1$. Let $\mathcal{B}_{s}:=\mathcal{B}((1-\epsilon)lh,\epsilon)$. From Lemma 9, we either find a vector $s\in\mathcal{B}_{s}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{s})\leq p2^{p(p-1)/4}$. Let $\mathcal{B}_{t}:=\mathcal{B}(-(1-\epsilon)lh,\epsilon)$. From Lemma 9, we either find a vector $t\in\mathcal{B}_{t}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{t})\leq p2^{p(p-1)/4}$.
Consider the subcase where Lemma 9 successfully found both vectors $s$ and $t$. Note that $\mathcal{B}_{s}$ and $\mathcal{B}_{t}$ are contained in $\mathcal{B}(0,1)$, thus $s,t$ are both feasible to Equation MITR. We apply Lemma 6 and obtain
$f(s)-f(t)\geq 2\left(1-5\epsilon\right)\lVert h\rVert_{2}\geq\left(1-7\epsilon\right)\lVert h\rVert_{2},$
where the last inequality holds for $\epsilon\in(0,1/3]$. Next, consider the subcase where Lemma 9 did not find both vectors $s$ and $t$. In this case, Lemma 9 found a vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B})\leq p2^{p(p-1)/4}$, for $\mathcal{B}\in\{\mathcal{B}_{s},\mathcal{B}_{t}\}$. We then have
$\operatorname{width}_{d}(\mathcal{B}(0,1))$ $=\frac{1}{\epsilon}\operatorname{width}_{d}(\mathcal{B})\leq\frac{1}{\epsilon}p2^{p(p-1)/4}.$
This concludes the proof in the second case. ∎
6 Flatness result for Equation E-MIQP
We are now ready to get back to Equation E-MIQP. The main goal of this section is to present a flatness result for this problem: an algorithm that either finds an $\epsilon$-approximate solution, or finds a nonzero vector $d$ along which the feasible region is flat. To state this result, and to work with ellipsoids, it will be useful to introduce the following notation. An ellipsoid is a set of the form
$\mathcal{E}(c,Q):=\left\{x\in\mathbb{R}^{n}:(x-c)^{\mathsf{T}}Q(x-c)\leq 1\right\},$
where $c\in\mathbb{R}^{n}$, and $Q\in\mathbb{R}^{n\times n}$ is a symmetric positive definite matrix. Note that $\mathcal{B}(c,r)=\mathcal{E}(c,I_{n}/r^{2})$, where $I_{n}$ is the $n\times n$ identity matrix. Notice that the inequality constraint in Equation E-MIQP can be equivalently written in the form $x\in\mathcal{E}(c,Q)$. We say that an ellipsoid $\mathcal{E}(c,Q)$ is rational when $c\in\mathbb{Q}^{n}$ and $Q\in\mathbb{Q}^{n\times n}$. We can now state the main result of this section.
Proposition 2 (Flatness result for Equation E-MIQP).
There is an algorithm which, for any $\epsilon\in(0,1]$, either finds an $\epsilon$-approximate solution to Equation E-MIQP, or finds a nonzero vector $d\in\mathbb{Z}^{n}$ with $d_{p+1}=\cdots=d_{n}=0$ and a scalar $\rho\in\mathbb{Q}$ such that
$\left\{d^{\mathsf{T}}x:x\in\mathcal{E}(c,Q)\right\}\subseteq\left[\rho,\rho+19p2^{p(p-1)/4}/\epsilon\right].$
The running time of the algorithm is polynomial in the size of Equation E-MIQP and in $\log(1/\epsilon)$.
The key ingredient in the proof of Proposition 2 is Proposition 1. However, we need three additional results, which are presented in Section 6.1.
6.1 Ingredients for flatness result
Given a rational ellipsoid $\mathcal{E}(c,Q)$, there may exist no rational affine transformation that maps $\mathcal{E}(c,Q)$ into the ball $\mathcal{B}(0,1)$. An example is given by the ellipsoid $\mathcal{E}(0,(2))$, as we mentioned in Section 1. The goal of the next lemma is to show that for any positive $\delta\in\mathbb{Q}$, we can find in polynomial time a rational affine transformation that maps $\mathcal{E}(c,Q)$ into an ellipsoid that is arbitrarily close to the ball $\mathcal{B}(0,1)$.
Lemma 10.
Let $\mathcal{E}(c,Q)\subset\mathbb{R}^{n}$ be a rational ellipsoid. There is an algorithm which, for any positive $\delta\in\mathbb{Q}$, finds a map $\tau:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\tau(x):=B(x-c)$, with $B\in\mathbb{Q}^{n\times n}$ invertible, such that
$\mathcal{B}(0,1)\subseteq\tau(\mathcal{E}(c,Q))\subseteq\mathcal{B}(0,1+\delta).$
The running time of the algorithm is polynomial in $\operatorname{size}(c),\operatorname{size}(Q),\log(1/\delta)$.
Proof.
From corollary 1 in [7], there is a strongly polynomial algorithm that finds an invertible matrix $M\in\mathbb{Q}^{n\times n}$ and a diagonal matrix $D\in\mathbb{Q}^{n\times n}$ such that $Q=M^{\mathsf{T}}DM$. Since $Q$ is positive definite, each diagonal element of $D$ is positive. We can then apply Lemma 8 to each diagonal element of $D$, and with $\epsilon:=\delta$. For $j=1,\dots,n$, we consider the element $d_{jj}$ of $D$ in position $(j,j)$, and we obtain a positive $l_{j}\in\mathbb{Q}$ such that $l_{j}\leq\sqrt{d_{jj}}\leq(1+\delta)l_{j}$. Let $L$ be the positive definite diagonal matrix in $\mathbb{Q}^{n\times n}$ with element $l_{j}$ in position $(j,j)$, for $j=1,\dots,n$.
Next, we show
$y^{\mathsf{T}}y\leq y^{\mathsf{T}}L^{-\mathsf{T}}DL^{-1}y\leq(1+\delta)^{2}y^{\mathsf{T}}y,\qquad\forall y\in\mathbb{R}^{n}.$ (4)
To see this, we use $d_{jj}\geq l_{j}^{2}$ for $j=1,\dots,n$, to obtain
$y^{\mathsf{T}}L^{-\mathsf{T}}DL^{-1}y=\sum_{j=1}^{n}\frac{d_{jj}}{l_{j}^{2}}y_{j}^{2}\geq\sum_{j=1}^{n}y_{j}^{2}=y^{\mathsf{T}}y,\qquad\forall y\in\mathbb{R}^{n},$
and we use $d_{jj}\leq(1+\delta)^{2}l_{j}^{2}$ for $j=1,\dots,n$, to derive
$y^{\mathsf{T}}L^{-\mathsf{T}}DL^{-1}y=\sum_{j=1}^{n}\frac{d_{jj}}{l_{j}^{2}}y_{j}^{2}\leq(1+\delta)^{2}\sum_{j=1}^{n}y_{j}^{2}=(1+\delta)^{2}y^{\mathsf{T}}y,\qquad\forall y\in\mathbb{R}^{n}.$
Let $\tau:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the map defined by $\tau(x):=(1+\delta)LM(x-c)$, thus with $B$ in the statement defined by $B:=(1+\delta)LM$. The containments in the statement then follow from (4) by writing the corresponding three sets as follows:
$\mathcal{B}(0,1)$ =$=\left\{y:y^{\mathsf{T}}y\leq 1\right\},$
$\tau(\mathcal{E}(c,Q))$ =$=\left\{\tau(x):(x-c)^{\mathsf{T}}Q(x-c)\leq 1\right\},$
=$=\left\{y:y^{\mathsf{T}}L^{-\mathsf{T}}M^{-\mathsf{T}}QM^{-1}L^{-1}y\leq(1+\delta)^{2}\right\},$
=$=\left\{y:y^{\mathsf{T}}L^{-\mathsf{T}}DL^{-1}y\leq(1+\delta)^{2}\right\},$
$\mathcal{B}(0,1+\delta)$ =$=\left\{y:y^{\mathsf{T}}y\leq(1+\delta)^{2}\right\}.$
The next result is a direct extension of Lemma 4, concerning the slightly different problem obtained from Equation TR by replacing the constraint $x\in\mathcal{B}(0,1)$ with $x\in\mathcal{B}(0,1+\delta)$:
$\begin{split}\min&\quad x^{\mathsf{T}}Hx+h^{\mathsf{T}}x\\ \operatorname{s.t.}&\quad x^{\mathsf{T}}x\leq(1+\delta)^{2}.\end{split}$ (TRδ)
Here $H$ is a symmetric matrix in $\mathbb{Q}^{n\times n}$, $h\in\mathbb{Q}^{n}$, and $\delta\in\mathbb{Q}$ is nonnegative.
Lemma 11.
Let $x^{\circ}$ be an optimal solution of Equation TRδ. There is an algorithm which, for any positive $\epsilon\in\mathbb{Q}$, finds a feasible solution $x^{\dagger}$ to Equation TRδ such that
$f(x^{\dagger})-f(x^{\circ})\leq\epsilon(1+\delta)^{2}\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}.$
The running time of the algorithm is polynomial in the size of Equation TRδ and in $\log(1/\epsilon)$.
Proof.
We apply the change of variables $y:=x/(1+\delta)$ to Equation TRδ and obtain
[…]
$f(x^{\dagger})-f(x^{\circ})$ $=f_{\delta}(y^{\dagger})-f_{\delta}(y^{\circ})$
$\leq\epsilon\max\left\{(1+\delta)^{2}\lVert H\rVert_{2},(1+\delta)\lVert h\rVert_{2}\right\}$
$\leq\epsilon(1+\delta)^{2}\max\left\{\lVert H\rVert_{2},\lVert h\rVert_{2}\right\}.$
The last result of the section discusses how flat directions in a mixed integer lattice are mapped to flat directions in the mixed integer lattice $\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$.
Lemma 12.
Let $b^{1},\dots,b^{n}$ be linearly independent vectors in $\mathbb{R}^{n}$ and let $p\in\{0,\dots,n\}$. Let $B\in\mathbb{R}^{n\times n}$ with columns $b^{1},\dots,b^{n}$, let $c\in\mathbb{R}^{n}$, and let $\tau:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the map defined by $\tau(x):=B(x-c)$. Let ${d^{\prime}}\in\mathbb{R}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$. Define $d:=B^{\mathsf{T}}{d^{\prime}}$. Then, $d\in\mathbb{Z}^{n}$, $d_{p+1}=\cdots=d_{n}=0$, and, for any nonempty bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$
Proof.
For $j=1,\dots,n$, we have $d_{j}={b^{j}}^{\mathsf{T}}d^{\prime}$. Thus, ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer implies $d_{1},\cdots,d_{p}$ integer, and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ implies $d_{p+1}=\cdots=d_{n}=0$. We now show that for any nonempty bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, we have $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$
$\operatorname{width}_{d}(\mathcal{S})$ $=\max\left\{d^{\mathsf{T}}x:x\in\mathcal{S}\right\}-\min\left\{d^{\mathsf{T}}x:x\in\mathcal{S}\right\}$
$=\max\left\{d^{\mathsf{T}}(x-c):x\in\mathcal{S}\right\}-\min\left\{d^{\mathsf{T}}(x-c):x\in\mathcal{S}\right\}$
$=\max\left\{{d^{\prime}}^{\mathsf{T}}B(x-c):x\in\mathcal{S}\right\}-\min\left\{{d^{\prime}}^{\mathsf{T}}B(x-c):x\in\mathcal{S}\right\}$
$=\max\left\{{d^{\prime}}^{\mathsf{T}}y:y\in\tau(\mathcal{S})\right\}-\min\left\{{d^{\prime}}^{\mathsf{T}}y:y\in\tau(\mathcal{S})\right\}$
$=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$
6.2 Proof of Proposition 2
We now have all ingredients that we need to prove Proposition 2. Before giving the proof, we give an overview of the algorithm. First, we construct an affine transformation $\tau$ that maps $\mathcal{E}(c,Q)$ into an ellipsoid that is sandwiched between $\mathcal{B}(0,1)$ and $\mathcal{B}(0,1+\epsilon)$. We denote by Equation E-MIQP the optimization problem obtained by applying the same affine transformation to Equation E-MIQP. Next, we consider Equation E-MIQP-w obtained from Equation E-MIQP by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with the weaker requirement $y\in\mathcal{B}(0,1+\epsilon)$ and by dropping the mixed integer lattice constraint. We then find an almost optimal solution $y^{\dagger}$ to Equation E-MIQP-w and use Lemma 9 to search for a close by vector $y^{\diamond}$ that is feasible to Equation E-MIQP. Next, we consider Equation E-MIQP-s obtained from Equation E-MIQP by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with with the stronger requirement $y\in\mathcal{B}(0,1)$. Using Proposition 1, we search for two feasible solutions $y^{1},y^{2}$ to Equation E-MIQP-s that are distant in objective value. If we manage to find all three vectors $y^{\diamond}$, $y^{1}$, $y^{2}$, we show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation E-MIQP. Otherwise, we show that $\mathcal{E}(c,Q)$ is flat and we find the associated vector $d$. We now give the complete proof of Proposition 2, which uses Lemmas 10, 11, 9, 1 and 12.
Proof.
Let $\epsilon^{\prime}:=\epsilon/18\in(0,1/18]$. Using the algorithm in Lemma 10 (with $\delta:=\epsilon^{\prime}$), we find a map $\tau:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\tau(x):=B(x-c)$, with $B\in\mathbb{Q}^{n\times n}$ invertible, such that
$\mathcal{B}(0,1)\subseteq\tau(\mathcal{E}(c,Q))\subseteq\mathcal{B}(0,1+\epsilon^{\prime}).$
Applying the change of variables $y:=B(x-c)$ to Equation E-MIQP, we obtain the optimization problem
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y\in\tau(\mathcal{E}(c,Q))\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\},\end{split}$ (E-MIQP)
where $\gamma:=c^{\mathsf{T}}Hc+{h}^{\mathsf{T}}c$, $H^{\prime}:=B^{-\mathsf{T}}HB^{-1}$, $h^{\prime}:=2B^{-\mathsf{T}}H^{\mathsf{T}}c+B^{-\mathsf{T}}h$, $b^{1},\dots,b^{n}$ are the columns of $B$, and $c^{\prime}:=-Bc$. For ease of notation, in the following we denote by $f_{y}(y):=y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$.
Consider the optimization problem obtained from Equation E-MIQP by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with the weaker requirement $y\in\mathcal{B}(0,1+\epsilon^{\prime})$ and by dropping the mixed integer lattice constraint:
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y^{\mathsf{T}}y\leq(1+\epsilon^{\prime})^{2}.\end{split}$ (E-MIQP-w)
Let $y^{\circ}$ be an optimal solution of Equation E-MIQP-w, which we do not need to compute. Using the algorithm in Lemma 11 (with $\epsilon:=\delta:=\epsilon^{\prime}$), we find a feasible solution $y^{\dagger}$ to Equation E-MIQP-w such that
$f_{y}(y^{\dagger})-f_{y}(y^{\circ})\leq\epsilon^{\prime}(1+\epsilon^{\prime})^{2}\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.$ (5)
Let $\mathcal{B}_{\diamond}:=\mathcal{B}((1-\epsilon^{\prime})y^{\dagger}/(1+\epsilon^{\prime}),\epsilon^{\prime})$. From Lemma 9, we either find a vector $y^{\diamond}\in\mathcal{B}_{\diamond}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\})$, or a nonzero vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d^{\prime}}(\mathcal{B}_{\diamond})\leq p2^{p(p-1)/4}$.
Consider now the mixed integer trust region problem obtained from Equation E-MIQP by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with the stronger requirement $y\in\mathcal{B}(0,1)$:
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y^{\mathsf{T}}y\leq 1\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\}.\end{split}$ (E-MIQP-s)
Using the algorithm in Proposition 1, we either find two feasible solutions $y^{1},y^{2}$ to Equation E-MIQP-s such that
$f(y^{1})-f(y^{2})\geq\left(1-7\epsilon^{\prime}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\},$ (6)
or we find a nonzero vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1))\leq p2^{p(p-1)/4}/\epsilon^{\prime}$.
In the remainder of the proof, we consider separately two cases. First, consider the case where Lemmas 9 and 1 successfully found all three vectors $y^{\diamond}$, $y^{1}$, $y^{2}$. Note that, since $y^{\dagger}\in\mathcal{B}(0,1+\epsilon^{\prime})$, we have $y^{\dagger}/(1+\epsilon^{\prime})\in\mathcal{B}(0,1)$, thus $\mathcal{B}_{\diamond}\subseteq\mathcal{B}(0,1)$ and $y^{\diamond}$ is feasible to Equation E-MIQP-s. Hence, the three vectors $y^{\diamond}$, $y^{1}$, $y^{2}$ are feasible to Equation E-MIQP-s and to Equation E-MIQP. Denote by $f_{\sup}$ and $f_{\inf}$ the supremum and the infimum of $f_{y}(y)$ on the feasible region of Equation E-MIQP. From (6), we directly obtain the following lower bound on $f_{\sup}-f_{\inf}$:
$\begin{split}f_{\sup}-f_{\inf}&\geq f(y^{1})-f(y^{2})\\ &\geq\left(1-7\epsilon^{\prime}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.\end{split}$ (7)
Next, we obtain an upper bound on $f_{y}(y^{\diamond})-f_{\inf}$. First, let $g\in\mathbb{R}^{n}$ so that $y^{\diamond}=y^{\dagger}+g$. It is simple to check that $\mathcal{B}_{\diamond}\subseteq\mathcal{B}(y^{\dagger},3\epsilon^{\prime})$, which implies $\lVert g\rVert_{2}\leq 3\epsilon^{\prime}$. Then,
$\lvert f_{y}(y^{\diamond})-f_{y}(y^{\dagger})\rvert$ $=\lvert f_{y}(y^{\diamond})-f_{y}(y^{\diamond}-g)\rvert$
$=\lvert{y^{\diamond}}^{\mathsf{T}}H^{\prime}y^{\diamond}+{h^{\prime}}^{\mathsf{T}}y^{\diamond}-(y^{\diamond}-g)^{\mathsf{T}}H^{\prime}(y^{\diamond}-g)-{h^{\prime}}^{\mathsf{T}}(y^{\diamond}-g)\rvert$
$=\lvert 2{y^{\diamond}}^{\mathsf{T}}H^{\prime}g-g^{\mathsf{T}}H^{\prime}g+{h^{\prime}}^{\mathsf{T}}g\rvert$
$\leq\lvert 2{y^{\diamond}}^{\mathsf{T}}H^{\prime}g\rvert+\lvert g^{\mathsf{T}}H^{\prime}g\rvert+\lvert{h^{\prime}}^{\mathsf{T}}g\rvert$
$\leq 2\lVert y^{\diamond}\rVert_{2}\lVert H^{\prime}\rVert_{2}\lVert g\rVert_{2}+\lVert g\rVert_{2}^{2}\lVert H^{\prime}\rVert_{2}+\lVert{h^{\prime}}\rVert_{2}\lVert g\rVert_{2}$
$\leq 6\epsilon^{\prime}\lVert H^{\prime}\rVert_{2}+9{\epsilon^{\prime}}^{2}\lVert H^{\prime}\rVert_{2}+3\epsilon^{\prime}\lVert{h^{\prime}}\rVert_{2}$
$\leq\left(9\epsilon^{\prime}+9{\epsilon^{\prime}}^{2}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.$
Here, in the third inequality we used $\lVert y^{\diamond}\rVert_{2}\leq 1$, $\lVert g\rVert_{2}\leq 3\epsilon^{\prime}$. Using the above relation and (5), we obtain the following upper bound on $f_{y}(y^{\diamond})-f_{\inf}$:
$\begin{split}f_{y}(y^{\diamond})-f_{\inf}&\leq f_{y}(y^{\diamond})-f_{y}(y^{\circ})\\ &\leq\lvert f_{y}(y^{\diamond})-f_{y}(y^{\dagger})\rvert+(f_{y}(y^{\dagger})-f_{y}(y^{\circ}))\\ &\leq\left(9\epsilon^{\prime}+9{\epsilon^{\prime}}^{2}+\epsilon^{\prime}(1+\epsilon^{\prime})^{2}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}\\ &=\left(10\epsilon^{\prime}+11\epsilon^{\prime 2}+\epsilon^{\prime 3}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.\end{split}$ (8)
We are now ready to show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation E-MIQP. We have
$f(y^{\diamond})-f_{\inf}$ $\leq(10\epsilon^{\prime}+11\epsilon^{\prime 2}+\epsilon^{\prime 3})\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}$
$\leq 18\epsilon^{\prime}(1-7\epsilon^{\prime})\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}$
$\leq 18\epsilon^{\prime}(f_{\sup}-f_{\inf})$
$=\epsilon(f_{\sup}-f_{\inf}).$
Here, in the first inequality we used (8), the second inequality holds for $\epsilon^{\prime}\in(0,1/18]$, and in the third inequality we used (7). It then follows that $x^{\diamond}:=B^{-1}y^{\diamond}+c$ is an $\epsilon$-approximate solution to Equation E-MIQP. This concludes the first case.
Next, consider the case where Lemma 9 did not find the vector $y^{\diamond}$, or Proposition 1 did not find the two vectors $y^{1}$, $y^{2}$. In this case, we first show that we have found a nonzero vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that
$\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1))\leq\frac{1}{\epsilon^{\prime}}p2^{p(p-1)/4}.$
If Proposition 1 did not find the two vectors $y^{1}$, $y^{2}$, then we already know it. Otherwise, if Lemma 9 did not find the vector $y^{\diamond}$, then we have found a nonzero vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d^{\prime}}(\mathcal{B}_{\diamond})\leq p2^{p(p-1)/4}$. We obtain
$\operatorname{width}_{d^{\prime}}\left(\mathcal{B}(0,1)\right)$ $=\frac{1}{\epsilon^{\prime}}\operatorname{width}_{d^{\prime}}\left(\mathcal{B}_{\diamond}\right)\leq\frac{1}{\epsilon^{\prime}}p2^{p(p-1)/4}.$
We can now upper bound $\operatorname{width}_{d^{\prime}}\left(\mathcal{B}(0,1+\epsilon^{\prime})\right)$ as follows:
$\operatorname{width}_{d^{\prime}}\left(\mathcal{B}(0,1+\epsilon^{\prime})\right)$ $=(1+\epsilon^{\prime})\operatorname{width}_{d^{\prime}}\left(\mathcal{B}(0,1)\right)$
$\leq\frac{(1+\epsilon^{\prime})}{\epsilon^{\prime}}p2^{p(p-1)/4}$
$=\left(1+\frac{18}{\epsilon}\right)p2^{p(p-1)/4}$
$\leq\frac{19}{\epsilon}p2^{p(p-1)/4}.$
Now let $d:=B^{\mathsf{T}}{d^{\prime}}\in\mathbb{Q}^{n}$. From Lemma 12, $d\in\mathbb{Z}^{n}$, $d_{p+1}=\cdots=d_{n}=0$, and, for any nonempty bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$ If we let $\mathcal{S}:=\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime}))$, where $\tau^{\leftarrow}$ denotes the inverse of $\tau$, we obtain
$\operatorname{width}_{d}(\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime})))=\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1+\epsilon^{\prime}))\leq\frac{19}{\epsilon}p2^{p(p-1)/4}.$
Let $\rho:=\min\{{d}^{\mathsf{T}}x:x\in\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime}))\}$, and note that we can calculate it as follows:
$\rho$ $=\min\{{d}^{\mathsf{T}}(B^{-1}y+c):y\in\mathcal{B}(0,1+\epsilon^{\prime})\}$
$=\min\{{d^{\prime}}^{\mathsf{T}}y:y\in\mathcal{B}(0,1+\epsilon^{\prime})\}+d^{\mathsf{T}}c$
$=-(1+\epsilon^{\prime})\lVert d^{\prime}\rVert_{2}+d^{\mathsf{T}}c$
$=-(1+\epsilon/18)\lVert d^{\prime}\rVert_{2}+d^{\mathsf{T}}c.$
We obtain
$\{d^{\mathsf{T}}x:x\in\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime}))\}\subseteq[\rho,\rho+19p2^{p(p-1)/4}/\epsilon].$
Since $\mathcal{E}(c,Q)\subseteq\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime}))$, we also have
$\{d^{\mathsf{T}}x:x\in\mathcal{E}(c,Q)\}\subseteq[\rho,\rho+19p2^{p(p-1)/4}/\epsilon].$
This concludes the second case. ∎
7 Approximation algorithm for Equation E-MIQP
In this section we prove Theorem 1. Before diving into the proof, we present, in Section 7.1, a few additional results that we need.
7.1 Ingredients for approximation algorithm
We start by presenting two lemmas that will be instrumental in reformulating a restricted Equation E-MIQP, where the feasible region is intersected with a hyperplane, as a new Equation E-MIQP with one fewer integer variable. In the first lemma we show that, given a hyperplane $\mathcal{H}$ in $\mathbb{R}^{n}$ that satisfies some technical conditions, we can find an affine transformation that maps the hyperplane $\{y\in\mathbb{R}^{n}:y_{1}=0\}$ to $\mathcal{H}$ and that preserves mixed integer points. Lemma 13 is similar to lemma 6 in [7]. The main difference is that here we seek an affine function from $\mathbb{R}^{n}$ to $\mathbb{R}^{n}$, while in lemma 6 in [7] the affine function is from $\mathbb{R}^{n-1}$ to $\mathcal{H}$. It will be useful to introduce some notation for hyperplanes. A hyperplane is a set of the form
$\mathcal{H}(d,\beta):=\{x\in\mathbb{R}^{n}:d^{\mathsf{T}}x=\beta\},$
where $d\in\mathbb{R}^{n}\setminus\{0\}$, $\beta\in\mathbb{R}$. We say that a hyperplane is rational, when $d\in\mathbb{Q}^{n}$ and $\beta\in\mathbb{Q}$. Recall that a square matrix $T\in\mathbb{Z}^{n\times n}$ is called unimodular if it has determinant ±$\pm 1$. The arguments used in the proof are direct extensions of those for pure integer linear programs (see, e.g., [4]).
Lemma 13.
Let $\mathcal{H}(d,\beta)\subset\mathbb{R}^{n}$ be a rational hyperplane with $d_{p+1}=\cdots=d_{n}=0$, for $p\in\{1,\dots,n\}$. There is a polynomial time algorithm which determines whether the set $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ is empty or not. If $\mathcal{S}\neq\emptyset$, the algorithm finds a map $\eta:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\eta(y):=\bar{x}+Ty$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T\in\mathbb{Z}^{n\times n}$ unimodular, such that
$\mathcal{H}(d,\beta)$ $=\eta(\{y\in\mathbb{R}^{n}:y_{1}=0\}),$
$\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$ $=\eta(\mathbb{Z}^{p}\times\mathbb{R}^{n-p}).$
Proof.
By possibly multiplying the equation $d^{\mathsf{T}}x=\beta$ by the least common multiple of the denominators of the entries of $d$, we may assume that $d$ is an integral vector. By possibly dividing the equation $d^{\mathsf{T}}x=\beta$ by the greatest common divisor of the entries of $d$, we may assume that $d$ has relatively prime entries. If $\beta\notin\mathbb{Z}$, then $\mathcal{S}$ is empty and we are done. Thus, we now assume $\beta\in\mathbb{Z}$. Since $d_{1},\dots,d_{p}$ are relatively prime, by corollary 1.9 in [4], the equation $\sum_{j=1}^{p}d_{j}x_{j}=\beta$ has an integral solution, thus $\mathcal{S}$ is nonempty. It is well known (see, e.g., proof of corollary 1.9 in [4]) that we can compute in polynomial time a vector $\tilde{x}\in\mathbb{Z}^{p}$ and a unimodular matrix $U\in\mathbb{Z}^{p\times p}$ such that
$\left\{x\in\mathbb{Z}^{p}:\sum_{j=1}^{p}d_{j}x_{j}=\beta\right\}=\{\tilde{x}+Uy:y\in\mathbb{Z}^{p},\ y_{1}=0\}.$
Taking the convex hull of the two sets in the above equality, we obtain
$\left\{x\in\mathbb{R}^{p}:\sum_{j=1}^{p}d_{j}x_{j}=\beta\right\}=\{\tilde{x}+Uy:y\in\mathbb{R}^{p},\ y_{1}=0\}.$
Since $\tilde{x}\in\mathbb{Z}^{p}$ and $U$ is unimodular we also have
$\mathbb{Z}^{p}=\{\tilde{x}+Uy:y\in\mathbb{Z}^{p}\}.$
We define the vector $\bar{x}\in\mathbb{Z}^{n}$ by $\bar{x}_{j}:=\tilde{x}_{j}$ for $j=1,\dots,p$, and $\bar{x}_{j}:=0$ for $j=p+1,\dots,n$. We also define the unimodular matrix $T\in\mathbb{Z}^{n\times n}$ with block corresponding to the first $p$ rows and $p$ columns being equal to $U$, block corresponding to the last $n-p$ rows and $n-p$ columns being equal to the identity matrix $I_{n-p}$, and remaining entries zero, i.e.,
$T:=\left(\begin{array}[]{c|c}U&0\\ \hline\cr 0&I_{n-p}\end{array}\right).$
Since $d_{i}=0$ for all $i\in\{p+1,\dots,n\}$, we conclude
$\mathcal{H}(d,\beta)$ =$=\{\bar{x}+Ty:y\in\mathbb{R}^{n},\ y_{1}=0\}$
$\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$ =$=\{\bar{x}+Ty:y\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}\}.$
The result then follows by defining the rational affine transformation $\eta:\mathbb{R}^{n}\to\mathbb{R}^{n}$, as $\eta(y):=\bar{x}+Ty$. ∎
The next lemma builds on Lemma 13 and is of key importance in our path to prove Theorem 1. Given a hyperplane $\mathcal{H}$ and an ellipsoid $\mathcal{E}$ in $\mathbb{R}^{n}$, we show that we can find an ellipsoid $\mathcal{E}^{\prime}$ in $\mathbb{R}^{n-1}$ and an affine transformation from $\mathbb{R}^{n-1}$ to $\mathbb{R}^{n}$, which maps $\mathcal{E}^{\prime}$ to $\mathcal{E}\cap\mathcal{H}$ and that preserves mixed integer points.
Lemma 14.
Let $\mathcal{E}(c,Q)\subset\mathbb{R}^{n}$ be a rational ellipsoid and let $\mathcal{H}(d,\beta)\subset\mathbb{R}^{n}$ be a rational hyperplane with $d_{p+1}=\cdots=d_{n}=0$, for $p\in\{1,\dots,n\}$. There is a polynomial time algorithm which determines whether the sets $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ and $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ are empty or not. If $\mathcal{S}\neq\emptyset$ and $n\geq 2$, the algorithm finds a map $\eta:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ of the form $\eta(y):=\bar{x}+Ty$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T\in\mathbb{Q}^{n\times{n-1}}$ of full rank, such that
$\mathcal{H}(d,\beta)$ $=\eta(\mathbb{R}^{n-1}),$
$\mathcal{S}$ $=\eta(\mathbb{Z}^{p-1}\times\mathbb{R}^{n-p}).$
If also $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)\neq\emptyset$, the algorithm finds the preimage of the set $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ under $\eta$, which is either a rational ellipsoid $\mathcal{E}(c^{\prime},Q^{\prime})\subset\mathbb{R}^{n-1}$ or a singleton $\left\{c^{\prime}\right\}$, for some $c^{\prime}\in\mathbb{Q}^{n-1}$.
Proof.
From Lemma 13, there is a polynomial time algorithm that determines whether the set $\mathcal{S}$ is empty or not. If $\mathcal{S}\neq\emptyset$, the algorithm finds a map $\eta^{\prime}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\eta^{\prime}(y):=\bar{x}+T^{\prime}y$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T^{\prime}\in\mathbb{Z}^{n\times n}$ unimodular, such that
$\mathcal{H}(d,\beta)$ $=\eta^{\prime}(\{y\in\mathbb{R}^{n}:y_{1}=0\})$
$\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$ $=\eta^{\prime}(\mathbb{Z}^{p}\times\mathbb{R}^{n-p}).$
The preimage of the ellipsoid $\mathcal{E}(c,Q)$ under $\eta^{\prime}$ is the ellipsoid
$\left\{y\in\mathbb{R}^{n}:(\bar{x}+T^{\prime}y-c)^{\mathsf{T}}Q(\bar{x}+T^{\prime}y-c)\leq 1\right\}=$
$=$ $\left\{y\in\mathbb{R}^{n}:(y-({T^{\prime}}^{-1}(c-\bar{x})))^{\mathsf{T}}{T^{\prime}}^{\mathsf{T}}Q{T^{\prime}}(y-({T^{\prime}}^{-1}(c-\bar{x})))\leq 1\right\}$
$=$ $\ \mathcal{E}({T^{\prime}}^{-1}(c-\bar{x}),{T^{\prime}}^{\mathsf{T}}Q{T^{\prime}}).$
For ease of notation, we define $q:={T^{\prime}}^{-1}(c-\bar{x})$ and $M:={T^{\prime}}^{\mathsf{T}}Q{T^{\prime}}$, so that the preimage of $\mathcal{E}(c,Q)$ under $\eta^{\prime}$ is $\mathcal{E}(q,M)$.
Let $T\in\mathbb{Q}^{n\times n-1}$ be the matrix $T^{\prime}$ stripped of the first column, and define the map $\eta:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ by $\eta(y):=\bar{x}+Ty$. We obtain $\mathcal{H}(d,\beta)=\eta(\mathbb{R}^{n-1})$ and $\mathcal{S}=\eta(\mathbb{Z}^{p-1}\times\mathbb{R}^{n-p}).$ Let $\mathcal{E}^{\prime}\subset\mathbb{R}^{n-1}$ be the preimage of the set $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ under $\eta$. We then have
:=$\mathcal{E}^{\prime}:=\left\{\tilde{y}\in\mathbb{R}^{n-1}:(0,\tilde{y})\in\mathcal{E}(q,M)\right\}.$
To conclude the proof, we need to show that $\mathcal{E}^{\prime}$ is either a rational ellipsoid $\mathcal{E}(c^{\prime},Q^{\prime})\subset\mathbb{R}^{n-1}$ or a singleton $\left\{c^{\prime}\right\}$, for some $c^{\prime}\in\mathbb{Q}^{n-1}$. Let $\tilde{q}$ be obtained from $q$ by dropping the first component. Define $m_{11}$ as the element of $M$ in position $(1,1)$, $\tilde{m}$ as the first column of $M$ without the first element, and $\tilde{M}$ is the submatrix of $M$ obtained by stripping $M$ of its first row and first column:
$q=\left(\begin{array}[]{c}q_{1}\\ \hline\cr\tilde{q}\end{array}\right),\quad M=\left(\begin{array}[]{c|c}m_{11}&\tilde{m}^{\mathsf{T}}\\ \hline\cr\tilde{m}&\tilde{M}\end{array}\right).$
We can now write $\mathcal{E}^{\prime}$ as follows:
$\mathcal{E}^{\prime}$ =$=\left\{\tilde{y}\in\mathbb{R}^{n-1}:(\tilde{y}-\tilde{q})^{\mathsf{T}}\tilde{M}(\tilde{y}-\tilde{q})-2q_{1}{\tilde{m}}^{\mathsf{T}}(\tilde{y}-\tilde{q})\leq 1-q_{1}^{2}m_{11}\right\}$
=$=\left\{\tilde{y}\in\mathbb{R}^{n-1}:\left(\tilde{y}-c^{\prime}\right)^{\mathsf{T}}\tilde{M}\left(\tilde{y}-c^{\prime}\right)\leq\zeta\right\},$
where $c^{\prime}:=\tilde{q}+q_{1}\tilde{M}^{-1}\tilde{m}$ and $\zeta:=1-q_{1}^{2}\left(m_{11}-\tilde{m}^{\mathsf{T}}\tilde{M}^{-1}\tilde{m}\right)$. Observe that $\mathcal{E}^{\prime}=\emptyset$ if $\zeta<0$, =$\mathcal{E}^{\prime}=\left\{c^{\prime}\right\}$ if $\zeta=0$, and $\mathcal{E}^{\prime}$ is the ellipsoid $\mathcal{E}(c^{\prime},\tilde{M}/\zeta)$ if $\zeta>0$. ∎
We remark that, in Lemma 14, it is fundamental that ellipsoids are encoded by means of their symmetric positive definite matrix, i.e., the matrix $Q$ in $\mathcal{E}(c,Q)$. A different encoding of the ellipsoid, where one encodes a matrix $L$ such that $L^{\mathsf{T}}L=Q$ instead of the matrix $Q$, would not allow us to obtain the ellipsoid $\mathcal{E}(c^{\prime},Q^{\prime})$ using only rational numbers.
We only need one more simple result, which illustrates a basic property of $\epsilon$-approximate solutions. The proof is given for completeness.
Lemma 15.
Let $\mathcal{S}^{1},\dots,\mathcal{S}^{t}\subseteq\mathbb{R}^{n}$. Let $f:\cup_{i=1}^{t}\mathcal{S}^{i}\to\mathbb{R}$ and assume that it has a (global) minimum. Let $\epsilon\in[0,1]$. For $i=1,\dots,t$, let $x^{i}$ be an $\epsilon$-approximate solution to the optimization problem $\inf\{f(x):x\in\mathcal{S}^{i}\}$. Then, each optimal solution to $\min\{f(x):x\in\{x^{1},\dots,x^{t}\}\}$, is an $\epsilon$-approximate solution to the optimization problem $\inf\{f(x):x\in\cup_{i=1}^{t}\mathcal{S}^{i}\}$.
Proof.
Let $f_{\inf}$ and $f_{\sup}$ denote the infimum and the supremum of $f(x)$ on the domain $\cup_{i=1}^{t}\mathcal{S}^{i}$. Let $x^{*}$ be an optimal solution to $\min\{f(x):x\in\cup_{i=1}^{t}\mathcal{S}^{i}\}$ and let $j\in\{1,\dots,t\}$ such that $x^{*}\in\mathcal{S}^{j}$. Let $f^{j}_{\inf}$ and $f^{j}_{\sup}$ denote the infimum and the supremum of $f(x)$ on the set $\mathcal{S}^{j}$. We then have
$f(x^{j})-f_{\inf}$ $=f(x^{j})-f^{j}_{\inf}$
$\leq\epsilon(f^{j}_{\sup}-f^{j}_{\inf})$
$\leq\epsilon(f_{\sup}-f_{\inf}).$
Here, in the first equality we used $f_{\inf}=f(x^{*})=f^{j}_{\inf}$, the first inequality holds because $x^{j}$ is an $\epsilon$-approximate solution to $\inf\{f(x):x\in\mathcal{S}^{j}\}$, and the last inequality holds because $f^{j}_{\sup}\leq f_{\sup}$ and $f^{j}_{\inf}=f_{\inf}$. Hence, $x^{j}$ is an $\epsilon$-approximate solution to $\inf\{f(x):x\in\cup_{i=1}^{t}\mathcal{S}^{i}\}$. Therefore, also each optimal solution to $\min\{f(x):x\in\{x^{1},\dots,x^{t}\}\}$ is an $\epsilon$-approximate solution to $\inf\{f(x):x\in\cup_{i=1}^{t}\mathcal{S}^{i}\}$. ∎
7.2 Proof of Theorem 1
We are now ready to prove Theorem 1. Before presenting the proof, we give an overview of the algorithm. First, we apply Proposition 2. If we find an $\epsilon$-approximate solution to Equation E-MIQP, we are done. Otherwise, we find a vector $d\in\mathbb{Z}^{n}$ that allows us to partition the feasible region of Equation E-MIQP into a number of parallel hyperplanes that is polynomial in $1/\epsilon$. Let $\mathcal{H}(d,\beta)$ be one of these hyperplanes and let Equation E-MIQP-$\beta$ be obtained from Equation E-MIQP by adding the constraint $d^{\mathsf{T}}x=\beta$. If $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ or $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ is empty, then Equation E-MIQP-$\beta$ is infeasible, so assume both sets are nonempty. If $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ is a singleton, then we can easily solve Equation E-MIQP-$\beta$. Otherwise, we use Lemma 14 to transform Equation E-MIQP-$\beta$ into an equivalent problem of the form of Equation E-MIQP, but with $p-1$ integer variables. We can then apply the algorithm recursively. We now give the complete proof of Theorem 1, which uses Propositions 2, 14 and 15.
Proof.
We now present our approximation algorithm for Equation E-MIQP.
Step 1: Approximation or partition. We first apply the algorithm in Proposition 2. If the algorithm finds an $\epsilon$-approximate solution to Equation E-MIQP, then we are done. Otherwise, the algorithm finds a nonzero vector $d\in\mathbb{Z}^{n}$ with $d_{p+1}=\cdots=d_{n}=0$ and a scalar $\rho\in\mathbb{Q}$ such that
$\{d^{\mathsf{T}}x:x\in\mathcal{E}(c,Q)\}\subseteq\left[\rho,\rho+19p2^{p(p-1)/4}/\epsilon\right].$
Note that $p\in\{1,\dots,n\}$ and $d^{\mathsf{T}}x$ is an integer for every $x\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$. Therefore, every feasible point of Equation E-MIQP is contained in one of the hyperplanes
$\{x\in\mathbb{R}^{n}:{d}^{\mathsf{T}}x=\beta\},\qquad\beta=\lceil\rho\rceil,\ \lceil\rho\rceil+1,\dots,\lfloor\rho+19p2^{p(p-1)/4}/\epsilon\rfloor.$
For each $\beta=\lceil\rho\rceil,\lceil\rho\rceil+1,\dots,\lfloor\rho+19p2^{p(p-1)/4}/\epsilon\rfloor$, we define the optimization problem
$\begin{split}\min&\quad x^{\mathsf{T}}Hx+h^{\mathsf{T}}x\\ \operatorname{s.t.}&\quad(x-c)^{\mathsf{T}}Q(x-c)\leq 1\\ &\quad d^{\mathsf{T}}x=\beta\\ &\quad x\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}.\end{split}$ (E-MIQP-$\beta$)
If $n=1$, Equation E-MIQP-$\beta$ can be trivially solved, thus we now assume $n\geq 2$. For each $\beta$, we apply the algorithm in Lemma 14. This algorithm determines whether the sets $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ and $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ are empty or not. If any of the two is empty, then Equation E-MIQP-$\beta$ is infeasible, thus we now assume that both sets are nonempty. Then, the algorithm finds a map $\eta:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ of the form $\eta(y):=\bar{x}+Ty$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T\in\mathbb{Q}^{n\times{n-1}}$ of full rank, and the preimage $\mathcal{E}^{\prime}$ of the set $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ under $\eta$. Assume first that $\mathcal{E}^{\prime}$ is a singleton $\left\{c^{\prime}\right\}$, for some $c^{\prime}\in\mathbb{Q}^{n-1}$. Then $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ is the singleton $\left\{\bar{x}+Tc^{\prime}\right\}$. We check if $\bar{x}+Tc^{\prime}\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$. If it is, then it is the optimal solution to Equation E-MIQP-$\beta$, and if it is not then Equation E-MIQP-$\beta$ is infeasible. Thus we now assume that $\mathcal{E}^{\prime}$ is an ellipsoid $\mathcal{E}(c^{\prime},Q^{\prime})\subset\mathbb{R}^{n-1}$. We then have
$\mathcal{H}(d,\beta)$ $=\eta(\mathbb{R}^{n-1}),$
$\mathcal{S}$ $=\eta(\mathbb{Z}^{p-1}\times\mathbb{R}^{n-p}),$
$\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ $=\eta(\mathcal{E}(c^{\prime},Q^{\prime})).$
Applying the change of variables $x=\bar{x}+Ty$ to Equation E-MIQP-$\beta$, we obtain the optimization problem
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y\in\mathcal{E}(c^{\prime},Q^{\prime})\\ &\quad x\in\mathbb{Z}^{p-1}\times\mathbb{R}^{n-p},\end{split}$ (MITR-$\beta^{\prime}$)
where $\gamma:=\bar{x}^{\mathsf{T}}H\bar{x}+{h}^{\mathsf{T}}\bar{x}$, $H^{\prime}:=T^{\mathsf{T}}HT$, $h^{\prime}:=T^{\mathsf{T}}{h}+2T^{\mathsf{T}}H^{\mathsf{T}}\bar{x}$. Note that Equation MITR-$\beta^{\prime}$ has the form of Equation E-MIQP, but with $p-1$ integer variables and $n-p$ continuous variables.
Step 2: Recursion. We apply recursively the algorithm to each obtained Equation MITR-$\beta^{\prime}$, for each $\beta=\lceil\rho\rceil,\lceil\rho\rceil+1,\dots,\lfloor\rho+19p2^{p(p-1)/4}/\epsilon\rfloor$. Since each time the problem is partitioned in Step 1, the number of integer variables decreases by one, the total number of iterations of the algorithm is upper bounded by
$\left(19p2^{p(p-1)/4}/\epsilon\right)^{p}.$
Step 3: Termination. If the algorithm did not find any feasible solution to Equation E-MIQP, the algorithm correctly returns that the feasible region is empty. Otherwise, among all the feasible solutions to Equation E-MIQP found, the algorithm returns one with the minimum objective value. The result then follows from Lemma 15. This is because the function $f(x):=x^{\mathsf{T}}Hx+h^{\mathsf{T}}x$, from $\mathcal{E}(c,Q)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ to $\mathbb{R}$, is a continuous function from a compact set to the real numbers. Hence, from the extreme value theorem, it has a (global) minimum. ∎
8 Mixed integer quadratic programming
In this section we prove Theorem 2.
8.1 Flatness result for Equation MIQP
Our first goal is to present a flatness result for Equation MIQP: an algorithm that either finds an approximate solution, or finds a nonzero vector $d$ along which the feasible region is flat. This result has several similarities with the flatness result for Equation E-MIQP that we presented in Proposition 2. The key difference, besides the optimization problem considered, is that the algorithm in Proposition 2 can be applied with any $\epsilon\in(0,1]$, while the algorithm in the flatness result for Equation MIQP only works with $\epsilon=1-\Theta(n^{-2})$.
Proposition 3 (Flatness result for Equation MIQP).
Consider Equation MIQP, and assume that $\{x\in\mathbb{R}^{n}:Wx\leq w\}$ is full-dimensional and bounded. There is an algorithm which either finds an $\epsilon$-approximate solution, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$, or finds a nonzero vector $d\in\mathbb{Z}^{n}$ with $d_{p+1}=\cdots=d_{n}=0$ and a scalar $\rho\in\mathbb{Q}$ such that
$\left\{d^{\mathsf{T}}x:Wx\leq w\right\}\subseteq\left[\rho,\rho+14(n+1)p2^{p(p-1)/4}\right].$
The running time of the algorithm is polynomial in the size of Equation MIQP.
Before giving the proof, we give an overview of the algorithm. The structure is similar to the algorithm in the proof of Proposition 2, with few key differences. First, we construct an affine transformation $\tau$ that maps $\mathcal{P}$ into a polytope that is sandwiched between $\mathcal{B}(0,1)$ and $\mathcal{B}(0,n+1)$. We denote by Equation MIQP the optimization problem obtained by applying the same affine transformation to Equation MIQP. Next, we consider Equation MIQP-s obtained from Equation MIQP by replacing the constraint $y\in\tau(\mathcal{P})$ with with the stronger requirement $y\in\mathcal{B}(0,1)$. Using Proposition 1, we search for two feasible solutions $y^{\#},y^{\diamond}$ to Equation MIQP-s such that $f(y^{\#})$ is sufficiently larger than $y^{\diamond}$. If Proposition 1 finds the vectors $y^{\#},y^{\diamond}$, we show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation MIQP, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$. Otherwise, we show that $\mathcal{P}$ is flat and we find the associated vector $d$. We now give the complete proof of Proposition 3, which uses Lemmas 10, 1 and 12.
Proof.
Let $\mathcal{P}:=\{x\in\mathbb{R}^{n}:Wx\leq w\}$. Using theorem 15.6 in [22] (with $\gamma:=1/(n+1/2)$), we find a rational ellipsoid $\mathcal{E}(c,Q)$ such that
$\mathcal{E}(c,(n+1/2)^{2}Q)\subseteq\mathcal{P}\subseteq\mathcal{E}(c,Q).$
Using the algorithm in Lemma 10 (with $Q:=(n+1/2)^{2}Q$ and $\delta:=1/(2n+1)$), we find a map $\tau:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\tau(x):=B(x-c)$, with $B\in\mathbb{Q}^{n\times n}$ invertible, such that
$\mathcal{B}(0,1)\subseteq\tau(\mathcal{E}(c,(n+1/2)^{2}Q))\subseteq\mathcal{B}(0,(2n+2)/(2n+1)).$
In particular, the first containment implies $\mathcal{B}(0,1)\subseteq\tau(\mathcal{P})$. Next, we prove $\tau(\mathcal{P})\subseteq\mathcal{B}(0,n+1)$:
$\tau(\mathcal{P})$ $\subseteq\tau(\mathcal{E}(c,Q))$
$=\mathcal{E}(0,B^{-\mathsf{T}}QB^{-1})$
$=(n+1/2)\ \mathcal{E}(0,(n+1/2)^{2}B^{-\mathsf{T}}QB^{-1})$
$=(n+1/2)\ \tau(\mathcal{E}(c,(n+1/2)^{2}Q))$
$\subseteq(n+1/2)\ \mathcal{B}(0,(2n+2)/(2n+1))$
$=\mathcal{B}(0,n+1).$
Therefore, we have shown
$\mathcal{B}(0,1)\subseteq\tau(\mathcal{P})\subseteq\mathcal{B}(0,n+1).$
Applying the change of variables $y:=B(x-c)$ to Equation MIQP, we obtain the optimization problem
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y\in\tau(\mathcal{P})\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\},\end{split}$ (MIQP)
where $\gamma:=c^{\mathsf{T}}Hc+{h}^{\mathsf{T}}c$, $H^{\prime}:=B^{-\mathsf{T}}HB^{-1}$, $h^{\prime}:=2B^{-\mathsf{T}}H^{\mathsf{T}}c+B^{-\mathsf{T}}h$, $b^{1},\dots,b^{n}$ are the columns of $B$, and $c^{\prime}:=-Bc$. For ease of notation, in the following we denote by $f_{y}(y):=y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$.
Consider now the mixed integer trust region problem obtained from Equation MIQP by replacing the constraint $y\in\tau(\mathcal{P})$ with the stronger requirement $y\in\mathcal{B}(0,1)$:
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y^{\mathsf{T}}y\leq 1\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\}.\end{split}$ (MIQP-s)
Using the algorithm in Proposition 1 (with $\epsilon:=1/14$), we either find two feasible solutions $y^{\#},y^{\diamond}$ to Equation MIQP-s such that
$f(y^{\#})-f(y^{\diamond})\geq\frac{1}{2}\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\},$ (9)
or we find a nonzero vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1))\leq 14p2^{p(p-1)/4}$.
In the remainder of the proof, we consider separately two cases. First, consider the case where Proposition 1 successfully found the two vectors $y^{\#},y^{\diamond}$. Note that $y^{\#}$ and $y^{\diamond}$ are feasible also to Equation MIQP. Denote by $f_{\sup}$ and $f_{\inf}$ the supremum and the infimum of $f_{y}(y)$ on the feasible region of Equation MIQP. From (9), we directly obtain the following lower bound on $f_{\sup}-f(y^{\diamond})$:
$\begin{split}f_{\sup}-f(y^{\diamond})&\geq f(y^{\#})-f(y^{\diamond})\\ &\geq\frac{1}{2}\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.\end{split}$ (10)
Next, we obtain the following upper bound on $f_{\sup}-f_{\inf}$:
$f_{\sup}-f_{\inf}\leq 4(n+1)^{2}\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.$ (11)
To prove this bound, observe that, for any $y\in\mathcal{B}(0,n+1)$,
$|f(y)|$ $\leq\lvert y^{\mathsf{T}}H^{\prime}y\rvert+\lvert{h^{\prime}}^{\mathsf{T}}y\rvert$
$\leq\lVert H^{\prime}\rVert_{2}\lVert y\rVert_{2}^{2}+\lVert h^{\prime}\rVert_{2}\lVert y\rVert_{2}$
$\leq(n+1)^{2}\lVert H^{\prime}\rVert_{2}+(n+1)\lVert h^{\prime}\rVert_{2}$
$\leq 2(n+1)^{2}\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.$
We are now ready to show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation MIQP, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$. We combine (10) and (11) to obtain:
$f_{\sup}-f_{\inf}\leq 4(n+1)^{2}\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}\leq 8(n+1)^{2}\left(f_{\sup}-f(y^{\diamond})\right).$
We can then use the above inequality to obtain the bound:
$f(y^{\diamond})-f_{\inf}$ $=(f_{\sup}-f_{\inf})-(f_{\sup}-f(y^{\diamond}))$
$\leq\left(1-\frac{1}{8(n+1)^{2}}\right)\left(f_{\sup}-f_{\inf}\right).$
It then follows that $x^{\diamond}:=B^{-1}y^{\diamond}+c$ is an $\epsilon$-approximate solution to Equation MIQP, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$. This concludes the first case.
Next, consider the case where Proposition 1 found a vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1))\leq 14p2^{p(p-1)/4}.$ We can now upper bound $\operatorname{width}_{d^{\prime}}\left(\mathcal{B}(0,n+1)\right)$ as follows:
$\operatorname{width}_{d^{\prime}}\left(\mathcal{B}(0,n+1)\right)$ $=(n+1)\operatorname{width}_{d^{\prime}}\left(\mathcal{B}(0,1)\right)$
$\leq 14(n+1)p2^{p(p-1)/4}.$
Now let $d:=B^{\mathsf{T}}{d^{\prime}}\in\mathbb{Q}^{n}$. From Lemma 12, $d\in\mathbb{Z}^{n}$, $d_{p+1}=\cdots=d_{n}=0$, and, for any nonempty bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$ If we let $\mathcal{S}:=\tau^{\leftarrow}(\mathcal{B}(0,n+1))$, where $\tau^{\leftarrow}$ denotes the inverse of $\tau$, we obtain
$\operatorname{width}_{d}(\tau^{\leftarrow}(\mathcal{B}(0,n+1)))=\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,n+1))\leq 14(n+1)p2^{p(p-1)/4}.$
Let $\rho:=\min\{{d}^{\mathsf{T}}x:x\in\tau^{\leftarrow}(\mathcal{B}(0,n+1))\}$, and note that we can calculate it as follows:
$\rho$ $=\min\{{d}^{\mathsf{T}}(B^{-1}y+c):y\in\mathcal{B}(0,n+1)\}$
$=\min\{{d^{\prime}}^{\mathsf{T}}y:y\in\mathcal{B}(0,n+1)\}+d^{\mathsf{T}}c$
$=-(n+1)\lVert d^{\prime}\rVert_{2}+d^{\mathsf{T}}c.$
We obtain
$\{d^{\mathsf{T}}x:x\in\tau^{\leftarrow}(\mathcal{B}(0,n+1))\}\subseteq[\rho,\rho+14(n+1)p2^{p(p-1)/4}].$
Since $\mathcal{P}\subseteq\tau^{\leftarrow}(\mathcal{B}(0,n+1))$, we also have
$\{d^{\mathsf{T}}x:x\in\mathcal{P}\}\subseteq[\rho,\rho+14(n+1)p2^{p(p-1)/4}].$
This concludes the second case. ∎
8.2 Approximation algorithm for Equation MIQP
We are now ready to prove Theorem 2. Before presenting the proof, we give an overview of the algorithm. First, we reduce ourselves to the case where the polyhedron $\left\{x\in\mathbb{R}^{n}:Wx\leq w\right\}$ is bounded and full-dimensional. Next, we apply Proposition 3. If we find an $\epsilon$-approximate solution to Equation MIQP where the approximation factor is $\epsilon=1-\Theta(n^{-2})$, then we are done. Otherwise, we find a vector $d\in\mathbb{Z}^{n}$ that allows us to partition the feasible region of Equation MIQP into a number of parallel hyperplanes that is polynomial in $n$. Let $\mathcal{H}(d,\beta)$ be one of these hyperplanes and let Equation MIQP-$\beta$ be obtained from Equation MIQP by adding the constraint $d^{\mathsf{T}}x=\beta$. We transform Equation MIQP-$\beta$ into an equivalent problem of the form of Equation MIQP, but with $p-1$ integer variables. We can then apply the algorithm recursively. We now give the complete proof of Theorem 2, which uses Propositions 3 and 15.
Proof.
We now present our approximation algorithm for Equation MIQP.
Step 0: Bounded feasible region. Consider Equation MIQP and assume there exists $\bar{f}\in\mathbb{R}$ such that every feasible solution has objective value at least $\bar{f}$. Theorem 4 in [9] implies that, if the feasible region in nonempty, there is an optimal solution of size bounded by an integer $\psi$, which is polynomial in the size of the problem.  [^1] Even though Theorem 4 in [9] does not give $\psi$ explicitly, a formula for $\psi$, as a function of the size of Equation MIQP, can be derived from its proof. We add to the system $Wx\leq w$ the linear inequalities $-2^{\psi}\leq x_{i}\leq 2^{\psi}$, for $i=1,\dots,n$, whose size is polynomial in the size of Equation MIQP. It is simple to check that an $\epsilon$-approximate solution to the obtained problem is also an $\epsilon$-approximate solution to the original, for every $\epsilon\in[0,1]$. Therefore, we can now assume that $\left\{x\in\mathbb{R}^{n}:Wx\leq w\right\}$ is bounded.
Step 1: Full dimensionality. Consider the sets
$\mathcal{P}:=\left\{x\in\mathbb{R}^{n}:Wx\leq w\right\},\qquad\mathcal{S}:=\mathcal{P}\cap\left(\mathbb{Z}^{p}\times\mathbb{R}^{n-p}\right),$
and note that $\mathcal{S}$ constitutes the feasible region of Equation MIQP. We apply the algorithm in theorem 1 in [8]. The algorithm either returns that $\mathcal{S}$ is empty, or it returns $p^{\prime}\in\left\{0,1,\dots,p\right\}$, $n^{\prime}\in\left\{p^{\prime},p^{\prime}+1,\dots,p^{\prime}+n-p\right\}$, a map $\tau:\mathbb{R}^{n^{\prime}}\to\mathbb{R}^{n}$ of the form $\tau(x^{\prime})=\bar{x}+Mx^{\prime}$, with $\bar{x}\in\mathbb{Z}^{p}\times\mathbb{Q}^{n-p}$ and $M\in\mathbb{Q}^{n\times n^{\prime}}$ of full rank, such that the polyhedron
$\mathcal{P}^{\prime}:=\left\{x^{\prime}\in\mathbb{R}^{n^{\prime}}:WMx^{\prime}\leq w-W\bar{x}\right\}$
is full-dimensional, and
$\mathcal{P}$ $=\tau\left(\mathcal{P}^{\prime}\right)$
$\mathcal{S}$ $=\tau\left(\mathcal{P}^{\prime}\cap\left(\mathbb{Z}^{p^{\prime}}\times\mathbb{R}^{n^{\prime}-p^{\prime}}\right)\right).$
If $\mathcal{S}$ is empty we are done. Otherwise, we perform the change of variables $x=\bar{x}+Mx^{\prime}$ to Equation MIQP and, after dropping the constant $\bar{x}^{\mathsf{T}}H\bar{x}+h^{\mathsf{T}}\bar{x}$ in the objective function, we obtain the problem
$\begin{split}\min&\quad{x^{\prime}}^{\mathsf{T}}H^{\prime}x^{\prime}+{h^{\prime}}^{\mathsf{T}}x^{\prime}\\ \operatorname{s.t.}&\quad W^{\prime}x^{\prime}\leq w^{\prime}\\ &\quad x^{\prime}\in\mathbb{Z}^{p^{\prime}}\times\mathbb{R}^{n^{\prime}-p^{\prime}},\end{split}$ (MIQP)
where $H^{\prime}:=M^{\mathsf{T}}HM$, ${h^{\prime}}^{\mathsf{T}}:=h^{\mathsf{T}}M+2\bar{x}^{\mathsf{T}}HM$, $W^{\prime}:=WM$, and $w^{\prime}:=w-W\bar{x}$. Since $\mathcal{P}$ is bounded, also the polyhedron $\mathcal{P}^{\prime}$, which is full-dimensional, is also bounded.
Since the definition of $\epsilon$-approximate solution is preserved under changes of variables and translations of the objective function, it then suffices to find an $\epsilon$-approximate solution to Equation MIQP. For ease of notation, from now on we consider Equation MIQP instead of Equation MIQP and we further assume, without loss of generality, that $\{x\in\mathbb{R}^{n}:Wx\leq w\}$ in Equation MIQP is full-dimensional.
Step 2: Approximation or partition. We now apply the algorithm in Proposition 3. If the algorithm finds an $\epsilon$-approximate solution to Equation MIQP where the approximation factor is $\epsilon=1-\Theta(n^{-2})$, then we are done. Otherwise, the algorithm finds a nonzero vector $d\in\mathbb{Z}^{n}$ with $d_{p+1}=\cdots=d_{n}=0$ and a scalar $\rho\in\mathbb{Q}$ such that
$\left\{d^{\mathsf{T}}x:Wx\leq w\right\}\subseteq\left[\rho,\rho+14(n+1)p2^{p(p-1)/4}\right].$
Note that $p\in\{1,\dots,n\}$ and $d^{\mathsf{T}}x$ is an integer for every $x\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$. Therefore, every feasible point of Equation MIQP is contained in one of the hyperplanes
$\{x\in\mathbb{R}^{n}:{d}^{\mathsf{T}}x=\beta\},\qquad\beta=\lceil\rho\rceil,\ \lceil\rho\rceil+1,\dots,\lfloor\rho+14(n+1)p2^{p(p-1)/4}\rfloor.$
For each $\beta=\lceil\rho\rceil,\lceil\rho\rceil+1,\dots,\lfloor\rho+14(n+1)p2^{p(p-1)/4}\rfloor$, we define the optimization problem
$\begin{split}\min&\quad x^{\mathsf{T}}Hx+h^{\mathsf{T}}x\\ \operatorname{s.t.}&\quad Wx\leq w\\ &\quad d^{\mathsf{T}}x=\beta\\ &\quad x\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}.\end{split}$ (MIQP-$\beta$)
If $n=1$, Equation MIQP-$\beta$ can be trivially solved, thus we now assume $n\geq 2$. For each $\beta$, we apply theorem 1 in [8] to the feasible region of Equation MIQP-$\beta$. If we find out that the feasible region is empty, we are done. Otherwise, we transform, with a change of variables, Equation MIQP-$\beta$ into a new problem, which we denote by Problem MIQP-$\beta^{\prime}$. Since $d_{p+1}=\cdots=d_{n}=0$, Problem MIQP-$\beta^{\prime}$ has at most $p-1$ integer variables and $n-p$ continuous variables.
Step 3: Recursion. We apply recursively the algorithm, from Step 1, to each obtained Problem MIQP-$\beta^{\prime}$, for each $\beta=\lceil\rho\rceil,\lceil\rho\rceil+1,\dots,\lfloor\rho+14(n+1)p2^{p(p-1)/4}\rfloor$. Since each time the problem is partitioned in Step 2, the number of integer variables decreases by one, the total number of iterations of the algorithm is upper bounded by
$\left(14(n+1)p2^{p(p-1)/4}\right)^{p}.$
Step 4: Termination. If the algorithm did not find any feasible solution to Equation MIQP, the algorithm correctly returns that the feasible region is empty. Otherwise, among all the feasible solutions to Equation MIQP found, the algorithm returns one with the minimum objective value. The result then follows from Lemma 15. This is because the function $f(x):=x^{\mathsf{T}}Hx+h^{\mathsf{T}}x$, from the original feasible region $\{x\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}:Wx\leq w\}$ to $\mathbb{R}$, is lower bounded by assumption, and thus has a (global) minimum, due to theorem 4 in [9]. ∎
Funding: A. Del Pia is partially funded by AFOSR grant FA9550-23-1-0433. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the Air Force Office of Scientific Research.
Acknowledgments: The author would like to thank two anonymous referees for comments and suggestions that improved the quality of this manuscript.